### v1

# Strategic Learning and Trading in Broker-Mediated Markets

Abstract
We study strategic interactions in a broker-mediated market. A broker provides liquidity to an informed trader and to noise traders while managing inventory in the lit market. The broker and the informed trader maximise their trading performance while filtering each other’s private information; the trader estimates the broker’s trading activity in the lit market while the broker estimates the informed trader’s private signal. Brokers hold a strategic advantage over traders who rely solely on prices to filter information. We find that information leakage in the client’s trading flow yields an economic value to the broker that is comparable to transaction costs; she speculates profitably and mitigates risk effectively, which, in turn, adversely impacts the informed trader’s performance. In contrast, low signal-to-noise sources, such as prices, result in the broker’s trading performance being indistinguishable from that of a naive strategy that internalises noise flow, externalises informed flow, and offloads inventory at a constant rate.
journal: TBA
1 Introduction
We study a model where a broker (she) provides liquidity to a strategic informed trader (he) and to zero-intelligence noise traders by acting as their counterparty. The broker charges transaction fees to both types of traders. Simultaneously, the broker manages her inventory exposure by trading in the lit market, where she faces both transaction costs and price impact. In this setting, the strategic agents (the broker and the informed trader) operate under partial and heterogeneous information. The informed trader observes a private signal, yet his trading strategy must account for the unobserved broker’s trading activity and its impact on prices, which he estimates from observed market prices. Meanwhile, the broker manages her exposure in the lit market without access to the private signal, which she estimates from her clients’ trading flow or from market price movements.
In this paper, we show how each strategic agent estimates their missing information by filtering relevant observed processes. Specifically, the informed trader relies solely on market prices, which, in a liquid market, have a low signal-to-noise ratio, resulting in low-quality estimates of the broker’s trading speed. In contrast, the broker has access to higher-quality information through her clients’ trading flow, as well as market prices. We show that, in practice, filtering based solely on price information does not produce accurate estimates of the private signal. However, the clients’ trading flow produces significantly better estimates, and can thus improve trading performance and protect the broker’s inventory from adverse selection. We find that the broker holds a strategic advantage over an informed trader who trades exclusively with her; thus, in practice, informed traders need to conceal information or distribute their trading flow across multiple brokers to maintain privacy. [^1] See, for example, Oomen, (2017), for a study of aggregators of multiple brokers.
We employ stochastic control tools in a linear-quadratic framework to derive the optimal adaptive strategies of both agents when they employ prices as the source of information. Both agents maximise their terminal expected wealth subject to inventory penalties. The performance criterion we introduce scales the risk aversion of both agents according to the conditional variance of their respective estimators. We characterise the optimal behaviour of both players and provide conditions for the existence and uniqueness of optimal Markovian strategies, which we derive explicitly. More precisely, given the linear-quadratic structure of the problems, we employ the separation principle and characterise the solutions to the control problems with systems of linear and matrix Riccati differential equations. The matrix Riccati differential equation that characterises the solution to the control problem of the broker is nonstandard. We show that the matrix Riccati differential equation admits a unique solution, which, in turn, ensures a unique optimal Markovian strategy for the broker, when the permanent price impact parameter lies within a certain interval $[0,\delta)$.
We run simulations to study the adaptive strategies of both agents. Our results show that market prices provide poor information for the filtering problems of the informed trader and the broker. Specifically, the broker’s price impact is insufficient to produce a reliable estimate of her trading activity, and the volatility of market prices makes it challenging to infer the informed trader’s private signal accurately. Consequently, the performance of the broker’s strategy in this scenario is akin to that of a naive strategy that externalises the toxic flow, internalises the noise flow, and unwinds inventory based on a trade-weighted average price schedule. In contrast, when the broker employs the trading flow of the informed trader to estimate the signal, her trading performance improves significantly. This improvement arises for two reasons: first, higher-quality signals enable the broker to speculate more effectively on future prices; second, she can manage externalisation and internalisation with better precision. Specifically, the broker increases externalisation and speculative activity when signal values are high or when she is more certain about her estimate.
Literature review
Our work combines elements of the execution literature for liquidity takers, and the market making literature for liquidity providers. Early works on algorithmic trading include Bertsimas and Lo, (1998), Ho and Stoll, (1983), Almgren and Chriss, (2001), and Avellaneda and Stoikov, (2008), which spurred numerous subsequent studies on execution and market making; see e.g., the textbooks Cartea et al., (2015) and Guéant, (2016). Our work also relates to stochastic games between strategic agents. Bank et al., (2021) develop a continuous-time trading game with two interacting markets: the dealer market where agents trade at an equilibrium price and liquidity providers charge a fee to execute liquidity takers’ demands, and the open market where the demands in the first market are absorbed in an unaffected, martingale price setting. Fu et al., (2022) introduce a portfolio liquidation game with self-exciting order flow that uses Hawkes processes with intensities that depend on past orders. Micheli et al., (2023) study open-loop, closed-loop, and central planning equilibrium of a game in which a interact through the impact they have in prices. Herdegen et al., (2023) investigate market makers competing for market orders of clients in a one-shot Nash competition. Abi Jaber et al., (2023) study a general class of infinite-dimensional linear-quadratic stochastic games with mean-field interaction between the agents. Chen et al., (2023) consider an optimal liquidation model where an investor executes meta-orders which trigger child orders and have both instantaneous and permanent feedback on future order flow. In our paper, the flow of information plays a key role. Here, only the broker has impact in prices because she is the only one that trades in the lit market (where trading is anonymous), and this impact is unknown to the informed trader. On the other hand, the informed trader observes a trading signal that the broker does not observe. Here, both agents (the broker and the informed trader) filter observed processes in search of the information they do not possess.
The combination of filtering techniques and stochastic control tools has been employed in the literature to study algorithmic trading. Campi and Zabaljauregui, (2020) study market making with filtering in a stochastic impulse control setting, Lim and Quenez, (2015) solve a portfolio optimisation problem under partial information for agents with no insider information in a stochastic price jump setting, Allan and Cohen, (2019) formulate a control problem to estimate parameters in the continuous-time Kalman-Bucy filter, Drissi, (2022) proposes models for optimal execution when the drift is estimated from prices, and Allan and Cohen, (2020) use pathwise stochastic control to construct filters that are robust to parameter uncertainty. Recently, Knochenhauer et al., (2024) combine the Kalman-Bucy filter with a linear-quadratic Gaussian control problem to study the problem of costly information acquisition. Pham, (2001) is concerned with mean-variance hedging when the drift of asset prices is partially observable and Pham and Quenez, (2001) search for the optimal portfolio in a stochastic volatility model where investors only observe prices. Aase et al., (2012) investigate the optimal trading of an insider using filtering theory. Firoozi et al., (2018) study a multi-agent (mean-field) game in which the agents attempt to filter a hidden process driving an observable common process. Casgrain and Jaimungal, (2018) characterise optimality in a large stochastic game with heterogeneous agents in the mean-field limit where agents filter latent factors. Finally, Cartea et al., 2024a employ filtering techniques to estimate the fundamental value of an asset from the trading flow of liquidity takers in decentralised trading venues. In this paper, the informed trader and the broker filter distinct information from each other. The broker does not observe the price signal while the informed trader does not observe the permanent price impact from the broker’s activity. As such, we introduce different filtering objects to both of them instead of a common filtering framework.
Closest to our work are Cartea and Sánchez-Betancourt, (2022) and Bergault and Sánchez-Betancourt, (2024). Both papers study versions of a game between a broker and her clients. [^2] There are other strands of the literature that are close to our work, such as Butz and Oomen, (2019) and Cartea et al., (2023) who develop tools to decide when to internalise or externalise in foreign exchange markets. Recently, Barzykin et al., (2024) study filtering of toxic flow in the externalisation of order flow. In the algorithmic trading literature, trading with signals appears in Cartea and Jaimungal, (2016) as the order flow of the market, or more generally in Neuman and Voß, (2022) as a predictable finite-variation process affecting asset prices. In Cartea and Sánchez-Betancourt, (2022), the broker faces an informed and an uninformed trader. The broker and the informed trader optimise their criteria, and they include ambiguity aversion to disentangle their interactions. [^3] Recently, Cartea et al., 2024b solve the perfect information Nash equilibrium of this game. Here we do not assume perfect information; instead, both agents filter the information they do not have. On the other hand, Bergault and Sánchez-Betancourt, (2024) study a large number of informed traders, each of which have a private and a common signal. They characterise, in the mean-field limit, the Nash equilibrium of the trading strategies. In our paper we depart from these approaches. Both the informed trader, and the broker, attempt to filter the missing information in their respective frameworks. More precisely, the broker filters the signal that only the informed trader has, and the informed trader filters the trading speed of the broker, which is only known to the broker and that has permanent impact on prices. As far as we are aware, this is the first time such a problem is studied in the algorithmic trading literature; more precisely, to the best of our knowledge this is the first time that liquidity providers and liquidity takers optimise their continuous-time stochastic control problems while filtering the information that they do not have.
The remainder of the paper proceeds as follows. Section 2 reviews the key elements of the model. Section 3 formulates the filtering problem of the informed trader and solves his stochastic control problem. Section 4 introduces the filtering problem of the broker and solves her stochastic control problem. Here, we prove that there is a range of values that the permanent price impact parameter can take, for which one can show existence and uniqueness of the optimal solution. Section 4.1.2 proposes two alternative filters for how the broker extracts an estimate of the signal that she does not observe. Section 5 carries out simulations to study the performance of the strategies we derive against three benchmarks. For given model parameters, we find evidence that the strategy that externalises the order flow of the informed trader, and internalises the order flow of the uninformed trader while offloading inventory, has a performance that is indistinguishable from the optimal strategy we derive. This is because in our setting, as well as in practice, the signal-to-noise ratio is low, and thus, the filtered signal is not informative. We show that the alternative filters we build have an economic value that is in the order of magnitude as transaction costs. We study the robustness of the best-performing estimate and find that it is robust to inventory misspecification for the earlier parts of the trading day, and it has a poor performance towards the end of the trading day.
2 The Model
In our model, a broker (she) provides liquidity to a strategic informed trader (he) and to uninformed noise traders. This section introduces the general competition setup between the two strategic players; the informed trader and the broker. Section 3 solves the optimal trading problem of the informed trader, and Section 4 solves the optimal trading problem of the broker.
Let $T>0$ be a fixed trading horizon and let ${\mathfrak{T}}=[0,T]$. We fix a filtered probability space $\left(\Omega,{\mathcal{F}},({\mathcal{F}}_{t})_{t\in{\mathfrak{T}}},\mathbb{P}\right)$ satisfying the usual conditions of right continuity and completeness. The probability space supports four standard Brownian motions $W^{S}$, $W^{\alpha}$, $W^{B}$, and $W^{U}$. The full-information filtration $({\mathcal{F}}_{t})_{t\in{\mathfrak{T}}}$ is that generated by the four Brownian motions. We consider two sub-filtrations in our model: (i) the filtration $\left({\mathcal{F}}^{I}_{t}\right)_{t\in{\mathfrak{T}}}$ of the informed trader, and (ii) the filtration $\left({\mathcal{F}}^{B}_{t}\right)_{t\in{\mathfrak{T}}}$ of the broker, which we define below.
The broker provides liquidity to the informed trader and to the uninformed trader by acting as counterparty to their trades throughout the trading window ${\mathfrak{T}}$. The informed trader and the uninformed trader carry out their trades at speeds $\eta$ and $\xi$ respectively. Positive values of the trading speed denote buying, and negative values denote selling. As remuneration for this service, the broker charges transaction costs to both traders. In our model, the transaction costs charged by the broker are linear in her clients’ trading speed and are modelled by the positive coefficients $\mathfrak{b}$ for the informed trader, and $\mathfrak{c}$ for the uninformed trader. [^4] If the broker charges the same fee to all clients, then we take $\mathfrak{b}=\mathfrak{c}$. In parallel to providing liquidity, the broker manages her exposure and inventory by trading in the open (or lit) market at a trading speed $\nu$. We assume that the broker incurs linear transaction costs modelled by the positive constant $\mathfrak{a}$, and that her trading activity in the market generates a linear permanent impact modelled by the coefficient $\mathfrak{p}>0$.
The midprice of the asset is driven by a stochastic signal $\alpha$ which is privately observed by the informed trader. Let $\left(S_{t}\right)_{t\in{\mathfrak{T}}}$ denote the process describing the midprice of the asset, whose dynamics are
$\mathrm{d}S_{t}=(\mathfrak{p}\,\nu_{t}+\alpha_{t})\,\mathrm{d}t+\sigma^{S}\mathrm{d}W_{t}^{S}\,,\qquad S_{0}\in\mathbb{R}^{+}\,,$ (2.1)
where $\sigma^{S}>0$ is the price volatility, and $W^{S}$ is a standard Brownian motion. The signal process $\left(\alpha_{t}\right)_{t\in{\mathfrak{T}}}$ follows the Ornstein–Uhlenbeck dynamics
$\mathrm{d}\alpha_{t}=-\kappa^{\alpha}\,\alpha_{t}\,\mathrm{d}t+\sigma^{\alpha}\,\mathrm{d}W^{\alpha}_{t}\,,\qquad\alpha_{0}\in\mathbb{R}\,,$ (2.2)
where $\kappa^{\alpha}\geq 0$ is the mean-reversion parameter, $\sigma^{\alpha}>0$ is the volatility of the signal, and $W^{\alpha}$ is a Brownian motion. The signal and price processes may share common stochastic factors, so we assume that $W^{\alpha}$ and $W^{S}$ are correlated and write $\langle W^{S},W^{\alpha}\rangle_{t}=\rho\,t$, where $\rho\in[-1,1]$. [^5] See Drissi, (2022) and Bergault et al., (2022) for models of optimal execution and market making with Ornstein–Uhlenbeck dynamics and predictive signals.
Let $\left(Q^{B}_{t}\right)_{t\in{\mathfrak{T}}}$ and $\left(X^{B}_{t}\right)_{t\in{\mathfrak{T}}}$ be the processes that describe the inventory and the cash of the broker, respectively, with dynamics
$\mathrm{d}Q^{B}_{t}$ $=\left(\nu_{t}-\eta_{t}-\xi_{t}\right)\mathrm{d}t\,,\qquad Q^{B}_{0}=q^{B}\in\mathbb{R}\,,$ (2.3)
$\mathrm{d}X^{B}_{t}$ $=-\left(S_{t}+\mathfrak{a}\,\nu_{t}\right)\nu_{t}\,\mathrm{d}t+\left(S_{t}+\mathfrak{b}\,\eta_{t}\right)\eta_{t}\,\mathrm{d}t+\left(S_{t}+\mathfrak{c}\,\xi_{t}\right)\xi_{t}\,\mathrm{d}t\,,\qquad X^{B}_{0}=0\,,$ (2.4)
where $\mathfrak{a},\mathfrak{b},\mathfrak{c}$ are the transaction costs coefficients described above. Similarly, the informed trader’s inventory process $\left(Q^{I}_{t}\right)_{t\in{\mathfrak{T}}}$ and cash process $\left(X^{I}_{t}\right)_{t\in{\mathfrak{T}}}$ follow the dynamics
$\mathrm{d}Q^{I}_{t}$ $=\eta_{t}\,\mathrm{d}t\,,\qquad Q^{I}_{0}=q^{I}\in\mathbb{R}\,,$ (2.5)
$\mathrm{d}X^{I}_{t}$ $=-\left(S_{t}+\mathfrak{b}\,\eta_{t}\right)\eta_{t}\,\mathrm{d}t\,,\qquad X^{I}_{0}=0\,.$ (2.6)
Below, we use the notation $P^{I,\eta}$ to emphasise that a process $P$ describes a property of the informed trader and is controlled through the process $\eta.$ Similarly, we use the notation $P^{B,\nu}$ for the broker.
In our model, the broker manages her inventory while providing liquidity to her clients, and the informed trader exploits his private signal. Both agents assume parametric models for their competitor’s behaviour, and employ the information they observe to filter their competitor’s private information. On the one hand, the broker observes the midprice $S$ of the asset, and given that she is counterparty to the trades of her clients, she also observes the trading speed of the informed trader $\eta$ and the trading speed of the uniformed trader $\xi$. In the main model, the broker filters the midprice minus her impact on prices to estimate the private signal $\alpha$ (which she does not observe). On the other hand, the informed trader observes his signal $\alpha$ and the midprice $S$, but he does not observe the trading speed of the broker. The informed trader assumes a parametric form for the broker’s strategy, which he uses to estimate the broker’s trading speed $\nu$. Both agents employ the Kalman-Bucy filter in their respective control problems. Sections 3 and 4 derive the optimal strategies of both agents.
3 The problem of the informed trader
This section derives the optimal adaptive strategy of the informed trader.
3.1   Learning equations
In our model, the informed trader does not directly observe the trading activity of the broker in the open market, that is, he does not observe $\nu$. However, he observes the signal $\alpha$ and he knows that the broker’s trading activity has a permanent impact. Thus, to infer the broker’s trading rate, he uses the process $\left(Y_{t}\right)_{t\in{\mathfrak{T}}}$ with dynamics
$\mathrm{d}Y_{t}=\mathrm{d}S_{t}-\alpha_{t}\,\mathrm{d}t=\mathfrak{p}\,\nu_{t}\,\mathrm{d}t+\sigma^{S}\mathrm{d}W_{t}^{S}\,.$ (3.1)
In practice, the informed trader has little a-priori knowledge on the behaviour of the broker. Here, he considers a probability measure $\mathbb{P}^{I}$ where the dynamics of the price follow (2.1) but with the following parametric form for $\nu$, given by
$\mathrm{d}\nu_{t}=-\theta^{B}\,\nu_{t}\,\mathrm{d}t+\sigma^{B}\mathrm{d}W_{t}^{B}\,,$ (3.2)
where $\theta^{B},\sigma^{B}>0$, $\nu_{0}$ is a square-integrable random variable measurable with respect to ${\mathcal{F}}^{I}_{0}$, and $W^{B}$ is a Brownian noise independent of $W^{S}$. The dynamics (3.2) are a key modelling assumption to derive the closed-form solution to the informed trader’s adaptive control problem, i.e., with filtering. The parameters $\theta^{B}$ and $\sigma^{B}$ encode what the informed trader believes about the dynamics of the trading speed of the broker.
[…]
$\mathrm{d}\hat{\nu}_{t}$ $=-\theta^{B}\,\hat{\nu}_{t}\,\mathrm{d}t+\frac{1}{(\sigma^{S})^{2}}\,\mathfrak{p}\,\mathbb{V}^{I}_{t}\,\left(\mathrm{d}Y_{t}-\mathfrak{p}\,\hat{\nu}_{t}\,\mathrm{d}t\right)\,,$ (3.3)
$\mathrm{d}\mathbb{V}^{I}_{t}$ $=\left[(\sigma^{B})^{2}-2\,\theta^{B}\,\mathbb{V}^{I}_{t}-\frac{\mathfrak{p}^{2}\,\left(\mathbb{V}^{I}_{t}\right)^{2}}{(\sigma^{S})^{2}}\right]\,\mathrm{d}t\,.$ (3.4)
For an initial distribution of $\nu_{0}$, the informed trader solves (3.3) to obtain the estimate $\hat{\nu}$ of $\nu$, with confidence intervals given by the variance $\mathbb{V}^{I}$ in (3.4). In particular, the dynamics of the variance in (3.4) are described by an ordinary differential equation (ODE) of the Riccati-type that admits a closed-form solution. For simplicity, in what follows, the informed trader assumes $\nu_{0}=0$ so he sets $\mathbb{V}^{I}_{0}=0$.
Equipped with the estimate $\hat{\nu}$ of the broker’s strategy in (3.3), the informed trader derives that under $(\Omega,{\mathcal{F}}^{I},({\mathcal{F}}^{I}_{t})_{t\in{\mathfrak{T}}},\mathbb{P}^{I})$, the midprice $S$ follows the dynamics
$\mathrm{d}S_{t}=$ $(\mathfrak{p}\,\hat{\nu}_{t}+\alpha_{t})\,\mathrm{d}t+\sigma^{S}\,\mathrm{d}\tilde{W}_{t}^{S,I}\,,$ (3.5)
where $\hat{\nu}$ is as in (3.3) and $\tilde{W}^{S,I}_{t}$ is a $(\mathbb{P}^{I},({\mathcal{F}}^{I}_{t})_{t\in{\mathfrak{T}}})$-Brownian motion.
Remark 1.
Note that the filtration $({\mathcal{F}}^{I}_{t})_{t\in{\mathfrak{T}}}$ is now fixed. This is because the latent $\nu$ is assumed to be exogenous and does not depend on the informed trader’s control. As such, the observation process $Y$ is already decoupled from the informed trader’s control and there is no “chicken and egg” problem as in Section 9.1 of Bensoussan, (2018).
3.2   The model
The informed trader exploits her trading signal to speculate and maximise her trading performance. However, he is averse to uncertainty about his competitor’s trading behaviour. More precisely, the set of admissible strategies is
[…]
and for every $\eta\in{\mathcal{A}}^{I}$, the informed trader maximises the performance criterion
$H^{I,\eta}(t,\mathfrak{x})=\mathbb{E}^{I}_{t,\mathfrak{x}}\left[X^{I,\eta}_{T}+Q^{I,\eta}_{T}\,S_{T}-(\beta_{0}+\beta_{1}\,\mathbb{V}^{I}_{T})\,\left(Q^{I,\eta}_{T}\right)^{2}-\int_{t}^{T}(\rho_{0}+\rho_{1}\,\mathbb{V}^{I}_{s})\left(Q^{I,\eta}_{s}\right)^{2}\,\mathrm{d}s\right]\,,$ (3.7)
where $\beta_{0},\beta_{1},\rho_{0},\rho_{1}>0$, $\mathfrak{x}=(s,x^{I},q^{I},\hat{\nu})$, and the expectation is under $\mathbb{P}^{I}$. The first two terms in the expectation in (3.7) are the informed trader’s terminal wealth composed of the marked-to-market value of his inventory and his terminal cash. The last two terms represent a terminal and running risk aversion components, respectively. In particular, we introduce the new terms $\beta_{1}\,\mathbb{V}^{I}$ and $\rho_{1}\,\mathbb{V}^{I}$ to adjust risk aversion according to the conditional variance $\mathbb{V}^{I}_{t}$. [^7] When $\beta_{1}=\rho_{1}=0$ we recover a performance criterion that is widely used in algorithmic trading; see e.g., Donnelly, (2022) for a recent review article. These terms penalise any terminal or running inventory according to the informed trader’s uncertainty because higher conditional variance implies more uncertainty in the estimation of the true $\nu$. The associated control problem is
$H^{I}(t,\mathfrak{x})=\sup_{\eta\in{\mathcal{A}}^{I}}H^{I,\eta}(t,\mathfrak{x})\,.$ (3.8)
3.3   Hamilton–Jacobi–Bellman equation
The Hamilton–Jacobi–Bellman (HJB) equation that characterises the value function $H^{I}$ is
00 0 $=\partial_{t}H^{I}-(\rho_{0}+\rho_{1}\,\mathbb{V}^{I}_{t})\left(q^{I}\right)^{2}+\sup_{\eta}\big{[}(\mathfrak{p}\,\hat{\nu}+\alpha)\,\partial_{s}H^{I}-\eta\left(s+\mathfrak{b}\,\eta\right)\,\partial_{x^{I}}H^{I}+\eta\,\partial_{q^{I}}H^{I}\big{]}$
$\qquad-\kappa^{\alpha}\alpha\,\partial_{\alpha}H^{I}-\theta^{B}\,\hat{\nu})\,\partial_{\hat{\nu}}H^{I}+\frac{1}{2}(\sigma^{S})^{2}\,\partial_{ss}H^{I}$
$\qquad+\mathfrak{p}\,\mathbb{V}^{I}_{t}\,\partial_{s\hat{\nu}}H^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\alpha}H^{I}+\frac{1}{2}\left(\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\right)^{2}\partial_{\hat{\nu}\hat{\nu}}H^{I}$ (3.9)
$\qquad+\sigma^{S}\,\sigma^{\alpha}\,\rho\,\partial_{s\alpha}H^{I}+\frac{\mathfrak{p}\,\sigma^{\alpha}\,\mathbb{V}^{I}_{t}\,\rho}{\sigma^{S}}\,\partial_{\alpha\hat{\nu}}H^{I}\,,$ (3.10)
with terminal condition
$H^{I}(T,\mathfrak{x})=x^{I}+q^{I}\,s-\left(\beta_{0}+\beta_{1}\,\mathbb{V}^{I}_{T}\right)\left(q^{I}\right)^{2}\,.$
The supremum in the Hamiltonian in (3.10) is attained at
$\eta^{*}=\frac{\partial_{q^{I}}H^{I}-\partial_{x^{I}}H^{I}\,s}{2\,\partial_{x^{I}}H^{I}\,\mathfrak{b}}\,.$ (3.11)
After substitution, employ the ansatz
$H^{I}=x^{I}+h^{I}(t,s,q^{I},\alpha,\hat{\nu})\,$
to obtain that $h^{I}$ satisfies the partial differential equation (PDE)
$0=-\left(\rho_{0}+\rho_{1}\,\mathbb{V}^{I}_{t}\right)\left(q^{I}\right)^{2}+\partial_{t}h^{I}+(\mathfrak{p}\hat{\nu}+\alpha)\,\partial_{s}h^{I}+\left(\partial_{q^{I}}h^{I}\right)^{2}/(4\,\mathfrak{b})-\kappa^{\alpha}\,\alpha\,\partial_{\alpha}h^{I}$
$\qquad-\theta^{B}\,\hat{\nu}\,\partial_{\hat{\nu}}h^{I}+\frac{1}{2}\left(\sigma^{S}\right)^{2}\partial_{ss}h^{I}+\mathfrak{p}\,\mathbb{V}^{I}_{t}\,\partial_{s\hat{\nu}}h^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\alpha}h^{I}$
$\qquad+\frac{1}{2}\left(\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\right)^{2}\partial_{\hat{\nu}\hat{\nu}}h^{I}+\frac{\mathfrak{p}\,\sigma^{\alpha}\,\mathbb{V}^{I}_{t}\,\rho}{\sigma^{S}}\,\partial_{\alpha\hat{\nu}}h^{I}\,,$
with terminal condition
$h^{I}(T,s,q^{I},\alpha,\hat{\nu})=q^{I}\,s-(\beta_{0}+\beta_{1}\,\mathbb{V}^{I}_{T})\left(q^{I}\right)^{2}\,.$
Next, use the ansatz
$h^{I}(t,s,q^{I},\alpha,\hat{\nu})=q^{I}\,s+g_{0}^{I}(t,\alpha,\hat{\nu})\,+q^{I}\,g_{1}^{I}(t,\alpha,\hat{\nu})+\left(q^{I}\right)^{2}\,g_{2}^{I}(t)$
to obtain the system of differential equations
00 0 $=-(\rho_{0}+\rho_{1}\mathbb{V}^{I}_{t})+(g_{2}^{I})^{\prime}+\frac{(g_{2}^{I})^{2}}{\mathfrak{b}}$ (3.12)
00 0 $=\partial_{t}\,g_{1}^{I}+(\mathfrak{p}\hat{\nu}+\alpha)+\frac{g_{1}^{I}\,g_{2}^{I}}{2\mathfrak{b}}-\kappa^{\alpha}\,\alpha\,\partial_{\alpha}\,g_{1}^{I}-\theta^{B}\,\hat{\nu}\,\partial_{\hat{\nu}}\,g_{1}^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\,\alpha}\,g_{1}^{I}\,,$
[…]
00 0 $=\partial_{t}\,g_{0}^{I}+\frac{(g_{1}^{I})^{2}}{4\mathfrak{b}}-\kappa^{\alpha}\,\alpha\,\partial_{\alpha}\,g_{0}^{I}-\theta^{B}\,\hat{\nu}\,\partial_{\hat{\nu}}\,g_{0}^{I}+\frac{1}{2}(\sigma^{\alpha})^{2}\partial_{\alpha\,\alpha}\,g_{0}^{I}$
[…]
with terminal conditions
$g_{2}^{I}(T)\quad$ $=-(\beta_{0}^{I}+\beta_{1}^{I}\,\mathbb{V}^{I}_{T})\,,$ (3.15)
$g_{1}^{I}(T,\alpha,\hat{\nu})$ $=\qquad 0\,,\qquad$ (3.16)
$g_{0}^{I}(T,\alpha,\hat{\nu})$ $=\qquad 0\,.\qquad$ (3.17)
The following lemma solves the system of PDEs (3.12)-(3.13)-(3.14).
Lemma 1.
There exist $g_{2}^{I},z_{i}^{I}\in C^{1}([0,T];\mathbb{R})$ for $i=0,\cdots,8$ such that $g_{2}^{I}$ and the following functions
$g_{1}^{I}(t,\alpha,\hat{\nu})$ $:=z_{1}^{I}(t)\,\alpha+z_{2}^{I}(t)\,\hat{\nu}\,,$ (3.18)
$g_{0}^{I}(t,\alpha,\hat{\nu})$ $:=z_{3}^{I}(t)+z_{4}(t)\,\alpha+z_{5}^{I}(t)\,\hat{\nu}+z_{6}^{I}(t)\,\alpha\,\hat{\nu}+z_{7}^{I}(t)\,\alpha^{2}+z_{8}^{I}(t)\,\hat{\nu}^{2}\,,$ (3.19)
solve the system of differential equations (3.14) with terminal conditions (3.15)-(3.16)-(3.17).
Proof.
Observe that (3.12) is a one-dimensional Riccati equation with quadratic term $0{1}/{\mathfrak{b}}>$ 0, “constant” term $-(\rho_{0}^{I}+\rho_{1}^{I}\,\mathbb{V}^{I}_{t})<0$, and terminal value $-(\beta_{0}^{I}+\beta_{1}^{I}\,\mathbb{V}^{I}_{t})<0$. Use Lemma 4.1 in Reid, (1972) to obtain that the equation has a unique solution in $(-\infty,T]$, and that it is negative.
Next, to solve (3.13) and (3.14), one may work backwards by substituting the desired form of $g_{0}$ and $g_{1}$ into the system of equations to obtain
00 0 $=\alpha$ $\left[1-\kappa^{\alpha}\,z_{1}^{I}+\frac{g_{2}^{I}\,z_{1}^{I}}{2\,\mathfrak{b}}+(z_{1}^{I})^{\prime}\right]$ (3.20)
[…]
00 0 $=$ $\left[\frac{\mathfrak{p}\,\sigma^{\alpha}\,\rho\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\,z_{6}^{I}+(\sigma^{\alpha})^{2}\,z_{7}^{I}+\left(\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}}{\sigma^{S}}\right)^{2}z_{8}^{I}+(z_{3}^{I})^{\prime}\right]$
[…]
Next, define a system of ODEs by setting all the terms in the square bracket above to be zero with $z_{i}(T)=0$ for $i=0,\cdots,8$. One solves the individual ODEs in the following order:
$(z_{1}^{I},z_{2}^{I})\to(z_{6}^{I},z_{7}^{I},z_{8}^{I})\to(z_{4}^{I},z_{5}^{I})\to z_{3}^{I}\,.$ (3.23)
In the above order, the ODEs are linear, so there exist $\{z_{i}^{I}\}_{i=1,\cdots,8}$ satisfying (3.13) and (3.14). ∎
3.4   The optimal adaptive strategy
The following theorem solves the adaptive control problem of the informed trader.
Theorem 1.
The HJB equation (3.10) admits a unique solution in $C^{1,2}([0,T],\mathbb{R}^{|\mathfrak{x}|})$ in the form of
$H^{I}(t,\mathfrak{x})=x^{I}+q^{I}\,s+g_{0}^{I}(t,\alpha,\hat{\nu})\,+q^{I}\,g_{1}^{I}(t,\alpha,\hat{\nu})+\left(q^{I}\right)^{2}\,g_{2}^{I}(t)\,,$ (3.24)
where $g_{i}^{I}$, $i=0,1,2$ are defined in Lemma 1. Moreover, there exists a constant $c<\infty$ such that
$|H^{I}(t,\mathfrak{x})|\leq c\,\left(1+\|\mathfrak{x}\|^{2}\right)\quad\forall(t,\mathfrak{x})\in[0,T]\times\mathbb{R}^{|\mathfrak{x}|}\,.$ (3.25)
Finally, the trader’s Markovian optimal trading rate for the control problem (3.8) in feedback form is
$\eta^{*}_{t}=\frac{z_{1}^{I}(t)\,\alpha_{t}+z_{2}^{I}(t)\,\hat{\nu}_{t}+2\,g_{2}^{I}(t)\,Q^{I*}_{t}}{2\,\mathfrak{b}}\,.$ (3.26)
where $z_{1}^{I},z_{2}^{I}\geq 0$ and $g_{2}^{I}<0$.
Proof.
Observe that (3.24) satisfies the HJB equation in (3.10). Moreover, by substituting $g_{i}^{I}$ for $i=0,1,2$ from Lemma 1, the solution to the HJB equation is a quadratic polynomial in terms of $\mathfrak{x}$ with time-dependent $C^{1}([0,T])$ coefficients; thus, the solution is bounded by a quadratic growth. By Theorem 3.5.2 (Verification Theorem) in Pham, (2009), the trader’s strategy (3.26) is indeed the optimal Markovian control. The last part of the theorem is as follows: $g_{2}^{I}$ is negative from Lemma 1.
Next, by solving for the ODE (3.21), we obtain
$z_{2}^{I}(t)=\int_{t}^{T}\mathfrak{p}\,\exp\left\{\int_{t}^{s}\frac{g_{2}^{I}(u)}{2\,\mathfrak{b}}-\theta^{B}\,\mathrm{d}u\right\}\,\mathrm{d}s\geq 0\quad\forall t\in[0,T]\,,$ (3.27)
which shows that $z_{2}^{I}$ is non-negative. The same argument follows for $z_{1}^{I}$ by solving (3.20). Observe that the equation above is linear in $\mathfrak{p}$. ∎
Remark 2.
If the informed trader were to observe $\nu$ and if $\nu$ was indeed an OU process, then the HJB equation of such full-information control problem would differ from (3.10) in the coefficients of $\partial_{\hat{\nu},\hat{\nu}}H^{I}$, $\partial_{s,\hat{\nu}}H^{I}$, and $\partial_{\alpha,\hat{\nu}}H^{I}$ only. From here, we use a version of Lemma 1 to show that the equivalent coefficients to $g_{2}^{I}$ and $g_{1}^{I}$ we obtain in such full-information case are exactly the same but changing $\hat{\nu}$ for $\nu$. This is the separation principle discussed in Section 9.3 of Bensoussan, (2018).
The optimal strategy (3.26) of the informed trader has three components. The first component is a speculative component that exploits the private signal. The second component uses the estimate that he has about the trading speed of the broker to speculate about future price movements due to the permanent price impact of the broker. The last component reduces the inventory of the informed trader according to his uncertainty and to the level of transaction costs charged by the broker.
4 The problem of the broker
This section derives the optimal adaptive strategy of the broker.
4.1   Learning equations
The broker assumes a model where the informed trader’s trading flow is driven by his private signal. More precisely, she assumes that the informed trader executes the strategy (3.26), so she assumes the parametric form
$2\,\mathfrak{b}\,\eta^{*}_{t}=z_{1}^{I}(t)\,\alpha_{t}+z_{2}^{I}(t)\,\hat{\nu}_{t}+2\,g_{2}^{I}(t)\,Q^{I*}_{t}\,,$ (4.1)
for deterministic functions $z_{i}(t)$ for $i=0,1,2$.
However, the broker does not observe $\alpha$, nor can she construct the filter $\hat{\nu}_{t}$ because the innovation process requires knowledge of $\alpha$. Thus, the broker replaces the signal $\alpha$ in (4.1) with an estimate $\hat{\alpha}$. Below, Section 4.1.1 assumes that the broker employs midprices to construct this estimate and in Section 4.1.2 the broker employs the trading speed of the informed trader as the source of information.
4.1.1 Learning from prices
Here, the broker assumes that her competitor’s trading speed takes the form
$\eta^{*}_{t}=\frac{z_{1}^{I}(t)\,\hat{\alpha}_{t}+z_{2}^{I}(t)\,\nu_{t}+2\,g_{2}^{I}(t)\,Q^{I*}_{t}}{2\,\mathfrak{b}}\,,$ (4.2)
where she replaces the estimate $\hat{\nu}$ with her true trading speed $\nu$ in (4.1)—this is equivalent to assuming that the filter $\hat{\nu}_{t}$ of the informed trader matches the true $\nu_{t}$. This latter assumption may be conservative, as perfect knowledge gives an obvious advantage to the informed trader. However, we show in the numerical examples of Section 5 that the broker can strategically influence (manipulate) the informed trader through this component, i.e., through knowing and exploiting that the informed trader’s strategy is based on that of the broker.
The broker observes the process
$\mathrm{d}Z_{t}=\mathrm{d}S_{t}-\mathfrak{p}\,\nu_{t}\,\mathrm{d}t=\alpha_{t}\,\mathrm{d}t+\sigma^{S}\,\mathrm{d}W^{S}_{t}\,.$ (4.3)
In (4.3), there is an implicit assumption that the trading speed of the broker is no longer an exogenously observed process, but instead it satisfies an equation that is linear in the filter $\hat{\alpha}$, the broker’s control $\nu$, and the integral of $\eta^{*}_{t}$ itself.
The broker projects $\alpha$ onto the filtration $\mathcal{Z}_{t}=\sigma(Z_{s}:\,s\in[0,t])\vee{\mathcal{N}}$, where ${\mathcal{N}}$ is the set of $\mathbb{P}$-null sets of the full sigma-algebra ${\mathcal{F}}$. Let $\hat{\alpha}_{t}=\mathbb{E}\left[\alpha_{t}\,|\,\mathcal{Z}_{t}\right]$ be the broker’s estimate and let $\mathbb{V}^{B}=\mathbb{E}\left[(\alpha_{t}-\hat{\alpha}_{t})^{2}\,|\,\mathcal{Z}_{t}\right]$ be its conditional variance. The Kalman Bucy filter gives
$\mathrm{d}\hat{\alpha}_{t}$ $=-\kappa^{\alpha}\,\hat{\alpha}_{t}\,\mathrm{d}t+\frac{1}{\left(\sigma^{S}\right)^{2}}\left(\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}\right)\,\left[\mathrm{d}Z_{t}-\hat{\alpha}_{t}\,\mathrm{d}t\right]\,,$ (4.4)
and
$\mathrm{d}\mathbb{V}^{B}_{t}$ $=\left[(1-\rho^{2})\,\left(\sigma^{\alpha}\right)^{2}+2\,\left(-\kappa^{\alpha}-\frac{1}{\sigma^{S}}\,\rho\,\sigma^{\alpha}\right)\,\mathbb{V}^{B}_{t}-\frac{1}{\left(\sigma^{S}\right)^{2}}\,\left(\mathbb{V}^{B}_{t}\right)^{2}\right]\,\mathrm{d}t\,.$ (4.5)
In particular, the conditional variance $\mathbb{V}^{B}$ of the estimator $\hat{\alpha}$ is a deterministic function of time. Finally, the broker fixes a probability measure $\mathbb{P}^{B}$ such that under $(\Omega,{\mathcal{F}}^{B},({\mathcal{F}}^{B}_{t})_{t\in{\mathfrak{T}}},\mathbb{P}^{B})$, the price follows the dynamics
$\mathrm{d}S_{t}$ $=(\mathfrak{p}\,\nu_{t}+\hat{\alpha}_{t})\,\mathrm{d}t+\sigma^{S}\,\mathrm{d}\tilde{W}_{t}^{S,B}\,,$ (4.6)
where $\tilde{W}_{t}^{S,B}$ is a $(\mathbb{P}^{B},({\mathcal{F}}^{B}_{t})_{t\in{\mathfrak{T}}})$-Brownian motion, and the filtration ${\mathcal{F}}^{B}_{t}=\sigma(S_{s},Z_{s},\xi_{s}:\,s\in[0,t])\vee{\mathcal{N}}$.
4.1.2 Learning from trading flow
The broker is counterparty to the trades of her clients; the informed and uninformed trader, for which she charges proportional transaction costs. Thus, she observes the trading flow of the informed trader. Here, we assume that the broker employs this trading flow as a source of information from which she obtains an estimate of the private signal $\alpha$. Here, $\eta^{*}$ is a measurable stochastic process that carries information about the unknown signal $\alpha$; c.f. with the assumptions in Section 4.1.1. The broker assumes the parametric form (4.1), so she observes the process ${\gamma}_{t}:=\eta^{*}_{t}-f_{3}(t)\,Q^{I*}_{t}=f_{1}(t)\,\alpha_{t}+f_{2}(t)\,\hat{\nu}_{t}$. We compute the dynamics of ${\gamma}_{t}$ to obtain
$\begin{split}\mathrm{d}{\gamma}_{t}&=\quad\left[f_{1}^{\prime}(t)\,\alpha_{t}-\kappa^{\alpha}\,f_{1}(t)\,\alpha_{t}+f_{2}^{\prime}(t)\,\hat{\nu}_{t}-\theta^{B}\,f_{2}(t)\,\hat{\nu}_{t}\right]\,\mathrm{d}t\\ &\qquad\qquad+\left[\sigma^{\alpha}\,f_{1}(t)\,\mathrm{d}W_{t}^{\alpha}+\frac{f_{2}(t)}{(\sigma^{S})^{2}}\,\mathfrak{p}\,\mathbb{V}^{I}_{t}\,\left(\mathrm{d}Y_{t}-\mathfrak{p}\,\hat{\nu}_{t}\,\mathrm{d}t\right)\right]\\ &=\quad\Bigg{[}\Big{(}f_{1}^{\prime}(t)-\kappa^{\alpha}\,f_{1}(t)\Big{)}\,\alpha_{t}+\left(f_{2}^{\prime}(t)-\theta^{B}\,f_{2}(t)-\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\left(\sigma^{S}\right)^{2}}\right)\,\hat{\nu}_{t}+\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\left(\sigma^{S}\right)^{2}}\,\nu_{t}\Bigg{]}\,\mathrm{d}t\\ &\qquad\qquad+\left[\sigma^{\alpha}\,f_{1}(t)\,\mathrm{d}W_{t}^{\alpha}+\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\sigma^{S}}\,\mathrm{d}W_{t}^{S}\right].\end{split}$ (4.7)
The broker knows that
$\hat{\nu}_{t}=\frac{\eta^{*}_{t}-f_{1}(t)\,\alpha_{t}-f_{3}(t)\,Q^{I*}_{t}}{f_{2}(t)}=\frac{\gamma_{t}-f_{1}(t)\,\alpha_{t}}{f_{2}(t)}\,,$
so she substitutes this in the above equation to obtain
$\begin{split}\mathrm{d}{\gamma}_{t}&=\left[\mathfrak{G}^{\alpha}(t)\,\alpha_{t}+\mathfrak{G}_{0}(t)\,{\gamma}_{t}+\mathfrak{G}_{1}(t)\,\nu_{t}+\mathfrak{G}_{2}(t)\right]\,\mathrm{d}t\\ &\qquad\qquad\qquad\qquad+\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]\,,\end{split}$ (4.8)
where, using (3.20)–(3.21), one obtains
$\mathfrak{G}_{0}(t)$ $=-\theta^{B}-\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}}{\left(\sigma^{S}\right)^{2}}+\frac{f_{2}^{\prime}(t)}{f_{2}(t)}$ (4.9)
$=-\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}}{\left(\sigma^{S}\right)^{2}}-\frac{\mathfrak{p}}{2\,f_{2}(t)\,\mathfrak{b}}-\frac{f_{3}(t)}{2}\,,$ (4.10)
$\mathfrak{G}^{\alpha}(t)$ $=f_{1}^{\prime}(t)-f_{1}(t)\,\left(\kappa^{\alpha}+\mathfrak{G}_{0}(t)\right)$ (4.11)
$=-\frac{1}{2\,\mathfrak{b}}-f_{1}(t)\,\left(\frac{f_{3}(t)}{2}+\mathfrak{G}_{0}(t)\right)\,,$ (4.12)
$\mathfrak{G}_{1}(t)$ $=\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\left(\sigma^{S}\right)^{2}},\quad\mathfrak{G}_{2}(t)=0\,,$ (4.13)
$\mathfrak{G}_{3}(t)$ $=\sigma^{\alpha}\,f_{1}(t),\quad\quad\mathfrak{G}_{4}(t)=\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\sigma^{S}}\,.$ (4.14)
We define the process $\tilde{Z}_{t}:=\gamma_{t}/\mathfrak{G}_{5}(t)$ with $\mathfrak{G}_{5}(t):=\sqrt{\mathfrak{G}_{3}(t)^{2}+\mathfrak{G}_{4}(t)^{2}+2\,\rho\,\mathfrak{G}_{3}(t)\,\mathfrak{G}_{4}(t)}$ that follows the dynamics
$\begin{split}\mathrm{d}\tilde{Z}_{t}&=\quad\left[\frac{-\tilde{Z}_{t}\,\mathfrak{G}_{5}^{{}^{\prime}}(t)}{\mathfrak{G}_{5}(t)}+\frac{\mathfrak{G}^{\alpha}(t)}{\mathfrak{G}_{5}(t)}\,\alpha_{t}+\frac{\mathfrak{G}_{0}(t)}{\mathfrak{G}_{5}(t)}\,\gamma_{t}+\frac{\mathfrak{G}_{1}(t)}{\mathfrak{G}_{5}(t)}\,\nu_{t}+\frac{\mathfrak{G}_{2}(t)}{\mathfrak{G}_{5}(t)}\right]\,\mathrm{d}t\\ &\qquad\qquad+\frac{1}{\mathfrak{G}_{5}(t)}\,\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]\,.\end{split}$ (4.15)
The coefficients in the above expression are well-defined when $t\to T$. Observe that
$\begin{split}\lim_{t\uparrow T}\tilde{Z}_{t}&=\lim_{t\uparrow T}\frac{f_{1}(t)}{\mathfrak{G}_{5}(t)}\,\alpha_{t}+\frac{f_{2}(t)}{\mathfrak{G}_{5}(t)}\,\hat{\nu}_{t}\\ &=\left(\lim_{t\uparrow T}\frac{f_{1}(t)}{\mathfrak{G}_{5}(t)}\right)\,\alpha_{T}+\left(\lim_{t\uparrow T}\frac{f_{2}(t)}{\mathfrak{G}_{5}(t)}\right)\,\hat{\nu}_{T}\,,\end{split}$ (4.16)
and from L’Hopital’s rule and the differential equations (3.20)–(3.21), it follows that $\lim_{t\uparrow T}{f_{1}(t)}/{\mathfrak{G}_{5}(t)}$ and $\lim_{t\uparrow T}{f_{2}(t)}/{\mathfrak{G}_{5}(t)}$ exist and are finite. This implies that $\tilde{Z}$ is a continuous square-integrable process. similarly, one also shows that the limits of ${\mathfrak{G}^{\alpha}(t)}/{\mathfrak{G}_{5}(t)}$, ${\mathfrak{G}_{1}(t)}/{\mathfrak{G}_{5}(t)}$, and ${\mathfrak{G}_{2}(t)}/{\mathfrak{G}_{5}(t)}$ exist and are finite as $t\uparrow T$.
On the other hand, observe that ${\mathfrak{G}_{5}^{{}^{\prime}}(t)}/{\mathfrak{G}_{5}(t)}$ and ${\mathfrak{G}_{0}(t)}/{\mathfrak{G}_{5}(t)}$ blow up as $t\uparrow T$. However, because $(\tilde{Z}_{t})_{t\in{\mathfrak{T}}}$ and $\big{(}\int_{0}^{t}\alpha_{s}\,{\mathfrak{G}^{\alpha}(s)}/{\mathfrak{G}_{5}(s)}+{\mathfrak{G}_{1}(s)}/{\mathfrak{G}_{5}(s)}\,\nu_{s}+{\mathfrak{G}_{2}(s)}/{\mathfrak{G}_{5}(s)}\,\mathrm{d}s\big{)}_{t\in{\mathfrak{T}}}$ are linear combinations of square-integrable progressively measurable continuous processes, the process $\big{(}\int_{0}^{t}{-\tilde{Z}_{s}\,\mathfrak{G}_{5}^{{}^{\prime}}(s)}/{\mathfrak{G}_{5}(s)}+\gamma_{s}\,{\mathfrak{G}_{0}(s)}/{\mathfrak{G}_{5}(s)}\,\mathrm{d}s\big{)}_{t\in{\mathfrak{T}}}$ is also square-integrable and progressively measurable. Next we define the main process that we use in the filtering framework of this section. The broker observes the process $Z_{t}$ defined as
$\begin{split}\mathrm{d}Z_{t}:&=\mathrm{d}\tilde{Z}_{t}-\mathfrak{f}_{t}\,\mathrm{d}t\\ &=\mathfrak{G}_{7}(t)\,\alpha_{t}\,\mathrm{d}t+\mathrm{d}\tilde{W}_{t}\,,\end{split}$ (4.17)
where $\mathfrak{f}_{t}=\mathfrak{G}_{6}(t)\,\tilde{Z}_{t}+\mathfrak{G}_{8}(t)\,\gamma_{t}+\mathfrak{G}_{9}(t)\,\nu_{t}+\mathfrak{G}_{10}(t)$ and
$\mathrm{d}\,\tilde{W}_{t}$ $:=\frac{1}{\mathfrak{G}_{5}(t)}\,\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]\,,$ (4.18)
$\mathfrak{G}_{6}(t)$ $=-\mathfrak{G}_{5}^{{}^{\prime}}(t)/\mathfrak{G}_{5}(t)\,,$ (4.19)
$\mathfrak{G}_{7}(t)$ $=\mathfrak{G}^{\alpha}(t)/\mathfrak{G}_{5}(t),\quad\mathfrak{G}_{8}(t)=\mathfrak{G}_{0}(t)/\mathfrak{G}_{5}(t)\,,$ (4.20)
$\mathfrak{G}_{9}(t)$ $=\mathfrak{G}_{1}(t)/\mathfrak{G}_{5}(t),\quad\mathfrak{G}_{10}(t)=\mathfrak{G}_{2}(t)/\mathfrak{G}_{5}(t)\,.$ (4.21)
Use the ODEs (3.20), (3.21), and (3.4), to rewrite $\mathfrak{G}_{6}(t)$ as
$\mathfrak{G}_{6}(t)=-\frac{\mathfrak{G}_{3}^{{}^{\prime}}(t)[\mathfrak{G}_{3}(t)+\rho\,\mathfrak{G}_{4}(t)]+\mathfrak{G}_{4}^{{}^{\prime}}(t)[\rho\,\mathfrak{G}_{3}(t)+\mathfrak{G}_{4}(t)]}{\mathfrak{G}_{5}(t)^{2}}\,,$ (4.22)
where
$\mathfrak{G}_{3}^{{}^{\prime}}(t)$ $=\sigma^{\alpha}\left[-\frac{1}{2\,\mathfrak{b}}+\kappa^{\alpha}\,f_{1}(t)-\frac{f_{3}(t)\,f_{1}(t)}{2}\right]\,,$ (4.23)
$\mathfrak{G}_{4}^{{}^{\prime}}(t)$ $=\frac{\mathfrak{p}}{\sigma^{S}}\left[\mathbb{V}^{I}_{t}\,\left(-\frac{\mathfrak{p}}{2\,\mathfrak{b}}+\theta^{B}\,f_{2}(t)-\frac{f_{3}(t)\,f_{2}(t)}{2}\right)+f_{2}(t)\,\left((\sigma^{B})^{2}-2\,\theta^{B}\,\mathbb{V}^{I}_{t}-\frac{\mathfrak{p}^{2}\,(\mathbb{V}^{I}_{t})^{2}}{(\sigma^{S})^{2}}\right)\right]\,.$ (4.24)
The broker sets up the filtering framework
$\mathrm{d}\alpha_{t}$ $=-\kappa^{\alpha}\,\alpha_{t}\,\mathrm{d}t+\sigma^{\alpha}\,\mathrm{d}W_{t}^{\alpha}\,,$ (4.25)
$\mathrm{d}Z_{t}$ $=\mathfrak{G}_{7}(t)\,\alpha_{t}\,\mathrm{d}t+\mathrm{d}\tilde{W}_{t}\,,$ (4.26)
where $Z$ is the observed process and $\alpha$ is the unobserved process. Finally, she takes $(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}}$ as the augmented filtration generated by the process $Z$. The proposition below computes, explicitly, the dynamics of the new filter.
Proposition 1.
Let $\hat{\alpha}^{\textup{alt}}_{t}:=\mathbb{E}\left[\alpha_{t}\,|\,\mathcal{Z}_{t}\right]$ and $\mathbb{V}^{\textup{alt},B}_{t}:=\mathbb{E}\left[(\alpha_{t}-\hat{\alpha}^{\textup{alt}}_{t})^{2}\,|\,\mathcal{Z}_{t}\right]$. The processes $\hat{\alpha}^{\textup{alt}}$ and $\mathbb{V}^{\textup{alt},B}$ satisfy the following stochastic differential equations,
$\mathrm{d}\hat{\alpha}^{\textup{alt}}_{t}$ $=-\kappa^{\alpha}\,\hat{\alpha}^{\textup{alt}}_{t}\,\mathrm{d}t+\left[\mathfrak{G}_{7}(t)\,\mathbb{V}^{\textup{alt},B}_{t}+\sigma^{\alpha}\,\mathfrak{K}(t)\right]\,\mathrm{d}I_{t}\,,$ (4.27)
$\mathrm{d}\mathbb{V}^{\textup{alt},B}_{t}$ $=\left[(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{V}^{\textup{alt},B}_{t}-(\mathfrak{G}_{7}(t)\,\mathbb{V}^{\textup{alt},B}_{t}+\sigma^{\alpha}\,\mathfrak{K}(t))^{2}\right]\,\mathrm{d}t\,,$ (4.28)
where $I$ is an innovation process satisfying the equation
$I_{t}:=Z_{t}-\int_{0}^{t}\mathbb{E}\left[\mathfrak{G}_{7}(s)\,\alpha_{s}\,\Big{|}\,\mathcal{Z}_{s}\right]\,\mathrm{d}s\,,$ (4.29)
and
$\mathfrak{K}(t)=\frac{\left(\mathfrak{G}_{3}(t)+\rho\,\mathfrak{G}_{4}(t)\right)}{\mathfrak{G}_{5}(t)}\,.$
The proof of Proposition 1, which is a slight extension of Corollary 3.38 in Bain and Crisan, (2009) for time-dependent coefficients, is in A.
4.2   The model
This section derives the closed-form optimal strategy of the broker when she employs the learning equations of Section 4.1.1, i.e., when she employs the midprice as the source of information. The derivation of the optimal adaptive strategy using the learning dynamics (4.28) of Section 4.1.2 is similar, and thus omitted. We remark that in the numerical experiments below, we study the performance of the broker’s strategy when she estimates the private signal according to (4.4)–(4.5) and also (4.27)–(4.28).
The broker assumes that the uninformed order flow $\xi$ is mean-reverting around zero and we write its dynamics as
$\mathrm{d}\xi_{t}$ $=-\kappa^{U}\,\xi_{t}\,\mathrm{d}t+\sigma^{U}\,\mathrm{d}W_{t}^{U}\,,\quad\xi_{0}=0\,.$ (4.30)
The inventory $Q^{B}$ of the broker evolves with the orders that she fills for her clients and her own trading in the lit market as follows
$\mathrm{d}Q^{B,\nu}$ $=\left(\nu_{t}-\eta^{*}_{t}-\xi_{t}\right)\,\mathrm{d}t\,,\quad Q^{B,\nu}_{0}=0\,,$ (4.31)
$\mathrm{d}Q^{I,\eta^{*}}_{t}$ $=\eta^{*}_{t}\,\mathrm{d}t\,,\quad Q^{I,\eta^{*}}_{0}=0\,.$ (4.32)
Similarly, the broker earns the transaction costs charged to her clients and incurs transaction costs from her own trading activity, so her cash $X^{B}$ follows the dynamics
$\mathrm{d}X^{B,\nu}_{t}$ =$=\left\{-\nu_{t}\,\left(S_{t}+\mathfrak{a}\,\nu_{t}\right)+\eta^{*}_{t}\,\left(S_{t}+\mathfrak{b}\,\eta^{*}_{t}\right)+\xi_{t}\,\left(S_{t}+\mathfrak{c}\,\xi_{t}\right)\right\}\,\mathrm{d}t\,,\quad X^{B,\nu}_{0}=0\,,$ (4.33)
where the transaction cost coefficients $\{\mathfrak{a},\mathfrak{b},\mathfrak{c}\}$ are described in Section 2.
The broker considers the following set of admissible controls for her trading strategy in the lit market:
[…]
For a given strategy $\nu\in{\mathcal{A}}^{B}$, she maximises the performance criterion
$H^{B,\nu}(t,\mathfrak{y})=\mathbb{E}_{t,\mathfrak{y}}\left[X^{B,\nu}_{T}+Q^{B,\nu}_{T}\,S_{T}-\left(\beta_{0}^{B}+\beta_{1}^{B}\,\mathbb{V}^{B}_{T}\right)\left(Q^{B,\nu}_{T}\right)^{2}-\int_{t}^{T}\left(\rho_{0}^{B}+\rho_{1}^{B}\,\mathbb{V}^{B}_{s}\right)\left(Q^{B,\nu}_{s}\right)^{2}\,\mathrm{d}s\right]$ (4.35)
where $\mathfrak{y}=(s,x^{B},q^{B},\hat{\alpha},\xi,q^{I})$, and the interpretation is similar to that of the informed trader. Thus, the control problem that the broker solves is
$H(t,\mathfrak{y})=\sup_{\nu\in{\mathcal{A}}^{B}}H^{B,\nu}(t,\mathfrak{y})\,.$ (4.36)
4.3   Hamilton–Jacobi–Bellman equation
The HJB equation associated with the value function $H$ in (4.36) is
[…]
$-\kappa^{\alpha}\,\hat{\alpha}\,\partial_{\hat{\alpha}}H+\frac{1}{2}\,\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}}{\sigma^{S}}\right)^{2}\,\partial_{\hat{\alpha},\hat{\alpha}}H+\left(\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}\right)\,\partial_{\hat{\alpha},s}H$ (4.38)
$-\nu\,\left(s+\mathfrak{a}\,\nu\right)\,\partial_{x^{B}}H+\eta^{*}\,\left(s+\mathfrak{b}\,\eta^{*}\right)\,\partial_{x^{B}}H+\xi\,\left(s+\mathfrak{c}\,\xi\right)\,\partial_{x^{B}}H$ (4.39)
$+\left(\nu-\eta^{*}-\xi\right)\,\partial_{q^{B}}H$ (4.40)
$+\left(\eta^{*}\right)\,\partial_{q^{I}}H-\kappa^{U}\,\xi\,\partial_{\xi}H+\frac{1}{2}\,\left(\sigma^{U}\right)^{2}\,\partial_{\xi\xi}H$ (4.41)
$-\left(\rho_{0}^{B}+\rho_{1}^{B}\,\mathbb{V}^{B}_{t}\right)\left(q^{B}\right)^{2}\bigg{\}}\,,$ (4.42)
with terminal condition
$H(T,\mathfrak{y})=x^{B}+q^{B}\,s-\left(\beta^{B}_{0}+\beta^{B}_{1}\,\mathbb{V}^{B}_{T}\right)\,\left(q^{B}\right)^{2}\,,$
where
$\eta^{*}=\frac{z_{1}^{I}(t)\,\hat{\alpha}+z_{2}^{I}(t)\,\nu+2\,g_{2}^{I}(t)\,q^{I}}{2\,\mathfrak{b}}\,.$ (4.43)
Next, define
$f_{1}(t)=z_{1}^{I}(t)/(2\,\mathfrak{b})\,,\quad f_{2}(t)=z_{2}^{I}(t)/(2\,\mathfrak{b})\,,\quad\text{and }f_{3}(t)=g_{2}^{I}(t)/\mathfrak{b}\,,$
and employ the ansatz
$H=x^{B}+h(t,s,q^{B},\hat{\alpha},\xi,q^{I})$
to obtain the following PDE in $h$
$\begin{split}0&=\partial_{t}h+\partial_{s}h\,\hat{\alpha}+\xi\left(s+\mathfrak{c}\,\xi\right)-\partial_{q^{B}}h\,\xi+\partial_{\hat{\alpha}}h\left(-\kappa^{\alpha}\hat{\alpha}\right)+\partial_{\xi}h\left(-\kappa^{U}\,\xi\right)\\ &\phantom{{}={}}+\frac{1}{2}\partial_{ss}h\left(\sigma^{S}\right)^{2}+\frac{1}{2}\partial_{\hat{\alpha}\hat{\alpha}}h\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\sigma^{\alpha}}{\sigma^{S}}\right)^{2}+\frac{1}{2}\partial_{\xi,\xi}h\left(\sigma^{U}\right)^{2}\\ &\phantom{{}={}}+\partial_{s\hat{\alpha}}h\left[\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}\right]-(\rho_{0}^{B}+\rho_{1}^{B}\,\mathbb{V}^{B}_{t})\,\left(q^{B}\right)^{2}\\ &\phantom{{}={}}+\left[\hat{\alpha}\,f_{1}+f_{3}\,q^{I}\right]\partial_{q^{I}}h-\left[\hat{\alpha}\,f_{1}+f_{3}\,q^{I}\right]\,\partial_{q^{B}}h\\ &\phantom{{}={}}+\Bigg{[}\hat{\alpha}^{2}\,f_{1}^{2}\,\mathfrak{b}+2\,\hat{\alpha}\,f_{1}\,f_{3}\,\mathfrak{b}\,q^{I}+f_{3}^{2}\,\mathfrak{b}\,\left(q^{I}\right)^{2}+\hat{\alpha}\,f_{1}\,s+f_{3}\,q^{I}\,s\Bigg{]}\\ &\phantom{{}={}}+\sup_{\nu}\Bigg{\{}-\left(\nu\right)^{2}\left(\mathfrak{a}-f_{2}^{2}\,\mathfrak{b}\right)+\nu\bigg{[}f_{2}\,\partial_{q^{I}}h+\left(1-f_{2}\right)\,\partial_{q^{B}}h\\ &\phantom{{}={}}\qquad+\left(2\,\hat{\alpha}\,f_{1}\,f_{2}\,\mathfrak{b}+2\,f_{2}\,f_{3}\,\mathfrak{b}\,q^{I}-s+f_{2}\,s\right)\,+\mathfrak{p}\,\partial_{s}h\bigg{]}\Bigg{\}}\,,\end{split}$ (4.44)
with terminal condition
$h(T,s,q^{B},\hat{\alpha},\xi,q^{I})=q^{B}\,s-\left(\beta^{B}_{0}+\beta^{B}_{1}\,\mathbb{V}^{B}_{T}\right)\,\left(q^{B}\right)^{2}\,.$
The supremum in (4.44) is attained at
$\nu^{*}=\frac{f_{2}\,\partial_{q^{I}}h+\left(1-f_{2}\right)\,\partial_{q^{B}}h+\left(2\,\hat{\alpha}\,f_{1}\,f_{2}\,\mathfrak{b}+2\,f_{2}\,f_{3}\,\mathfrak{b}\,q^{I}-s+f_{2}\,s\right)+\mathfrak{p}\,\partial_{s}h}{2\,\left(\mathfrak{a}-f_{2}^{2}\,\mathfrak{b}\right)}\,,$ (4.45)
because $\left(\mathfrak{a}-f_{2}^{2}\,\mathfrak{b}\right)>0$, so the function $h$ satisfies the PDE
$\begin{split}0&=\partial_{t}h+\partial_{s}h\,\hat{\alpha}+\xi\,\left(s+\mathfrak{c}\,\xi\right)-\partial_{q^{B}}h\,\xi+\partial_{\hat{\alpha}}h\left[-\kappa^{\alpha}\,\hat{\alpha}\right]+\partial_{\xi}h\left[-\kappa^{U}\,\xi\right]\\ &\phantom{{}={}}+\frac{1}{2}\partial_{ss}h\left(\sigma^{S}\right)^{2}+\frac{1}{2}\partial_{\hat{\alpha}\hat{\alpha}}h\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}}{\sigma^{S}}\right)^{2}+\frac{1}{2}\partial_{\xi,\xi}h\left(\sigma^{U}\right)^{2}\\ &\phantom{{}={}}+\partial_{s\hat{\alpha}}h\left[\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\sigma^{\alpha}\right]-(\rho_{0}^{B}+\rho_{1}^{B}\,\mathbb{V}^{B}_{t})\,q^{2}\\ &\phantom{{}={}}+\left[\hat{\alpha}\,f_{1}+f_{3}\,q^{I}\right]\partial_{q^{I}}h-\left[\hat{\alpha}\,f_{1}+f_{3}\,q^{I}\right]\,\partial_{q^{B}}h\\ &\phantom{{}={}}+\Bigg{[}\hat{\alpha}^{2}\,f_{1}^{2}\,\mathfrak{b}+2\,\hat{\alpha}\,f_{1}\,f_{3}\,\mathfrak{b}\,q^{I}+f_{3}^{2}\,\mathfrak{b}\,\left(q^{I}\right)^{2}+\hat{\alpha}\,f_{1}\,s+f_{3}\,q^{I}\,s\Bigg{]}\\ &\phantom{{}={}}+\frac{\left[f_{2}\,\partial_{q^{I}}h+\left(1-f_{2}\right)\,\partial_{q^{B}}h+\left(2\,\hat{\alpha}\,f_{1}\,f_{2}\,\mathfrak{b}+2\,f_{2}\,f_{3}\,\mathfrak{b}\,q^{I}-s+f_{2}\,s\right)+\mathfrak{p}\,\partial_{s}h\right]^{2}}{4\,\left(\mathfrak{a}-f_{2}^{2}\,\mathfrak{b}\right)}\,.\end{split}$ (4.46)
The HJB equation in matrix form
To study the problem further, we show that the PDE (4.46) is a system of matrix differential equations involving a matrix Riccati ODE. Employ the ansatz
$h(t,s,\tilde{\mathfrak{y}})=q^{B}\,s+G_{0}+2\,G_{1}\,\tilde{\mathfrak{y}}^{\top}+\tilde{\mathfrak{y}}\,G_{2}\,\tilde{\mathfrak{y}}^{\top}$ (4.47)
where $\tilde{\mathfrak{y}}=\begin{pmatrix}q^{B}&\hat{\alpha}&\xi&q^{I}\end{pmatrix}$, $G_{0}:[0,T]\to\mathbb{R}$, $G_{1}:[0,T]\to\mathbb{R}^{1\times 4}$, and $G_{2}:[0,T]\to\mathbb{R}_{\text{symmetric}}^{4\times 4}$ are functions of time, and write the HJB equation (4.46) as
00 0 $=$ $\Bigg{[}G_{0}^{\prime}+4\,G_{1}\,P_{8}^{\top}\,P_{8}\,G_{1}^{\top}+\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}}{\sigma^{S}}\right)^{2}\,e_{2}\,G_{2}\,e_{2}^{\top}+(\sigma^{U})^{2}\,e_{3}\,G_{2}\,e_{3}^{\top}\Bigg{]}$ (4.48)
[…]
where
$P_{2}$ $=\left(\begin{array}[]{ccccc}0&0&0&0\\ -{f_{1}}&-{\kappa^{\alpha}}&0&{f_{1}}\\ -1&0&-{\kappa^{U}}&0\\ -{f_{3}}&0&0&{f_{3}}\\ \end{array}\right)\,,$ (4.55)
$P_{5}$ $=\left(\begin{array}[]{ccccc}-{\mathbb{V}^{B}}{\rho_{1}^{B}}-{\rho_{0}^{B}}&1/2&0&0\\ 1/2&{f_{1}}^{2}{\mathfrak{b}}&0&{f_{1}}{f_{3}}{\mathfrak{b}}\\ 0&0&{\mathfrak{c}}&0&\\ 0&{f_{1}}{f_{3}}{\mathfrak{b}}&0&{f_{3}}^{2}{\mathfrak{b}}&\\ \end{array}\right)\,,$ (4.60)
$P_{7}$ $=\left(\begin{array}[]{ccccc}\frac{\mathfrak{p}}{2\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}&\frac{{f_{1}}{f_{2}}{\mathfrak{b}}}{\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}&0&\frac{{f_{2}}{f_{3}}{\mathfrak{b}}}{\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}\\ \end{array}\right)\,,$ (4.62)
$P_{8}$ $=\left(\begin{array}[]{ccccc}\frac{1-{f_{2}}}{2\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}&0&0&\frac{{f_{2}}}{2\sqrt{{\mathfrak{a}}-{f_{2}}^{2}{\mathfrak{b}}}}\\ \end{array}\right)\,,$ (4.64)
and $\{e^{i}\}$ is the canonical basis of $\mathbb{R}^{4}$.
Recall that $\tilde{\mathfrak{y}}\,M\,\tilde{\mathfrak{y}}^{\top}=\tilde{\mathfrak{y}}\,M^{\top}\,\tilde{\mathfrak{y}}^{\top}$ for any four-by-four matrix $M$ and define
$P_{9}=2\,P_{8}^{\top}\,P_{7}+P_{2}^{\top}\,,$ (4.65)
to obtain the matrix differential equations
00 0 $=$ $\,G_{0}^{\prime}+4\,G_{1}\,P_{8}^{\top}\,P_{8}\,G_{1}^{\top}+\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}}{\sigma^{S}}\right)^{2}\,e_{2}\,G_{2}\,e_{2}^{\top}+(\sigma^{U})^{2}\,e_{3}\,G_{2}\,e_{3}^{\top}\,,$ (4.66)
00 0 $=$ $\,G_{1}^{\prime}+G_{1}\,P_{2}^{\top}+2\,G_{1}\,P_{8}^{\top}\,P_{7}+4\,G_{1}\,P_{8}^{\top}\,P_{8}\,G_{2}\,,$ (4.67)
00 0 $=$ $\,G_{2}^{\prime}+P_{7}^{\top}\,P_{7}+4\,G_{2}^{\top}\,P_{8}^{\top}\,P_{8}\,G_{2}+G_{2}^{\top}\,P_{9}+P_{9}^{\top}\,G_{2}+P_{5}\,,$ (4.68)
with terminal conditions
$G_{0}=0\,,\quad G_{1}=0\,,$
and
$G_{2}(T)=\begin{pmatrix}-(\beta_{0}^{B}+\beta_{1}^{B}\,\mathbb{V}^{B}_{T})&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}\,.$ (4.69)
The next lemma shows existence and uniqueness of a solution to the differential equation in $G_{2}$. Here, and in the remainder of the paper we assume that the values of the model parameters are such that $01+\mathfrak{b}\,f_{3}(t)>$ 0 for all $t\in[0,T]$. Note that $\mathfrak{b}\,f_{3}/2=g_{2}^{I}$, thus, the inequality can be written as $g_{2}^{I}>-1/2$. Given $g_{2}^{I}$ satisfies the ODE in (3.12), we see that if the value of the temporary impact and terminal inventory risk parameters for the informed trader ($\mathfrak{b}$, $\beta_{0}^{I}$, and $\beta_{1}^{I}$) are small, this condition is met. In the numerical examples below the values of the model parameters are such that this condition is satisfied.
Proposition 2.
There exists $\delta>0$ such that for all $\mathfrak{p}\in[0,\delta)$, the matrix Riccati equation (4.68) admits a unique solution $G_{2}:[0,T]\to\mathbb{R}^{4\times 4}$ with terminal condition (4.69).
Proof.
The matrix Riccati equation in (4.68) can be simplified because the components of $P_{8}^{\top}\,P_{8}$ are zero except in the corners, and the first and last columns of $P_{9}$ take the form $\begin{pmatrix}\cdots&0&0&\cdots\end{pmatrix}^{\top}$. The simplified differential equation satisfied by $\tilde{G}_{2}$ is
$\begin{split}0=\tilde{G}_{2}^{\prime}+4\,\tilde{G}_{2}^{\top}\,U\,\tilde{G}_{2}+\tilde{G}_{2}^{\top}\,V+V^{\top}\,\tilde{G}_{2}+B\,,\quad\tilde{G}_{2}(T)=\begin{pmatrix}-(\beta_{0}^{B}+\beta_{1}^{B}\,\mathbb{V}^{B}_{T})&0\\ 0&0\end{pmatrix}\,,\end{split}$ (4.70)
with
$\begin{split}\tilde{G}_{2}&=\begin{pmatrix}(G_{2})_{11}&(G_{2})_{14}\\ (G_{2})_{14}&(G_{2})_{44}\end{pmatrix}\,,\\ U&=\frac{1}{\mathfrak{a}-f_{2}^{2}\mathfrak{b}}\,\begin{pmatrix}(1-f_{2})^{2}&f_{2}\,(1-f_{2})\\ f_{2}\,(1-f_{2})&f_{2}^{2}\end{pmatrix}=\begin{pmatrix}1-f_{2}\\ f_{2}\end{pmatrix}\,\left(\mathfrak{a}-f_{2}^{2}\mathfrak{b}\right)^{-1}\,\begin{pmatrix}1-f_{2}&f_{2}\end{pmatrix}\,,\\ V&=\frac{1}{\mathfrak{a}-f_{2}^{2}\mathfrak{b}}\,\begin{pmatrix}\frac{\mathfrak{p}\,(1-f_{2})}{2}&-f_{3}\,(\mathfrak{a}-f_{2}\,\mathfrak{b})\\ \frac{\mathfrak{p}\,f_{2}}{2}&\mathfrak{a}\,f_{3}\end{pmatrix}\,,\\ B&=\frac{1}{\mathfrak{a}-f_{2}^{2}\mathfrak{b}}\,\begin{pmatrix}\frac{\mathfrak{p}^{2}}{4}-\left(\mathfrak{a}-f_{2}^{2}\,\mathfrak{b}\right)\left(\rho_{0}^{B}+\rho_{1}^{B}\,\mathbb{V}^{B}_{t}\right)&\frac{\mathfrak{p}\,f_{2}\,f_{3}\,\mathfrak{b}}{2}\\ \frac{\mathfrak{p}\,f_{2}\,f_{3}\,\mathfrak{b}}{2}&f_{3}^{2}\,\mathfrak{a}\,\mathfrak{b}\end{pmatrix}\,.\end{split}$ (4.71)
Recall that $z_{2}^{I}$ satisfies (3.27), so $f_{2}(t)=\mathfrak{p}\,\int_{t}^{T}\exp\left\{\int_{t}^{s}\frac{f_{3}(u)}{2}-\theta^{B}\,\mathrm{d}u\right\}\,\mathrm{d}s$, $f_{2}=0$ whenever $\mathfrak{p}=0$, and $f_{2}$ is continuous in $\mathfrak{p}$. Next, use Theorem 2.3 in Freiling et al., (2000) by setting
$C=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}\,,\qquad\displaystyle D=-\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\,,$
and consider the following matrices:
$B_{11}$ $=V\,,\quad B_{12}=U\,,\quad B_{21}=-B^{\top}\,,\quad B_{22}=-V^{\top}\,,$ (4.72)
$E$ $=C+D\,\tilde{G}_{2}(T)+\left(D\,\tilde{G}_{2}(T)\right)^{\top}\,,$ (4.73)
$L(t)$ $=\begin{pmatrix}C\,B_{11}+D\,B_{21}&C\,B_{12}+B_{11}^{\top}\,D+D\,B_{22}\\ {0}&B_{12}^{\top}\,D\end{pmatrix}\,$ (4.74)
$=\begin{pmatrix}C\,V+B^{\top}&C\,U\\ {0}&-U^{\top}\end{pmatrix}\,.$ (4.75)
It follows that
$E=\begin{pmatrix}2\,(\beta_{0}^{B}+\beta_{1}^{B}\,\mathbb{V}^{B}_{t})&0\\ 0&1\end{pmatrix}>0\,,\quad\text{and}\quad L+L^{\top}=\begin{pmatrix}C\,V+V^{\top}\,C+2\,B&C\,U\\ U\,C&-2\,U\end{pmatrix}\,.$ (4.76)
[…]
Lastly, observe that the differential equation for $G_{1}$ is a factor of itself and the terminal condition is $\textbf{0}\in\mathbb{R}^{1\times 4}$. Thus, the vector-valued solution is $\textbf{0}^{1\times 4}$. The ODE satisfied by $G_{0}$ is linear so it has a unique solution.
4.4   The optimal adaptive strategy
The following theorem solves the adaptive control problem of the broker.
Theorem 2.
Let $\mathfrak{p}\in[0,\delta)$ with $\delta$ as in Proposition 2. Then, the HJB equation (4.42) admits a unique solution in $C^{1,2}([0,T],\mathbb{R}^{|\mathfrak{y}|})$ with quadratic growth in $\mathfrak{y}$. Moreover, the broker’s Markovian optimal trading rate for the control problem (4.36) in feedback form is given by
$\nu^{*}_{t}=\frac{1}{\sqrt{\mathfrak{a}-\mathfrak{b}\,f_{2}^{2}(t)}}\,\left(P_{7}+2\,P_{8}\,G_{2}(t)\right)\,\tilde{\mathfrak{y}}^{\top}_{t}\,,$ (4.77)
where $\tilde{\mathfrak{y}}_{t}=\begin{pmatrix}Q^{B*}_{t}&\hat{\alpha}_{t}&\xi_{t}&Q^{I}_{t}\end{pmatrix}$.
Proof.
By Proposition 2 there is a unique solution to the matrix Riccati differential equation (4.68). Armed with $G_{2}$, we find the unique solution $G_{0}$ to the linear differential equation (4.66). With $G_{0}$ and $G_{2}$, we obtain the following solution to the HJB equation (4.42):
$H^{B}(t,\mathfrak{y})=x^{B}+q^{B}\,s+G_{0}(t)+\tilde{\mathfrak{y}}\,G_{2}(t)\,\tilde{\mathfrak{y}}^{\top}\,.$ (4.78)
We observe that $H^{B}$ is a quadratic polynomial in $\mathfrak{y}$ with time-dependent $C^{1}([0,T],\mathbb{R})$ coefficients; thus, it is bounded by a quadratic growth. By Theorem 3.5.2 in Pham, (2009), the broker’s strategy (4.77) is indeed her optimal Markovian control. ∎
The optimal adaptive strategy (4.77) of the broker balances internalisation and externalisation of the trading flow of her clients and has four components. The first component reduces the inventory of the broker according to her model parameters—trading costs, trading revenue from clients, and aversion parameters. The second component exploits the estimated private signal of the informed trader. The third component adjust the strategy according to the trading activity of the uninformed traders. Finally, the last component adjusts the strategy according to the informed trader inventory level. Below, Section 5 thoroughly studies the internalisation and externalisation components in the strategy of the broker.
5 Numerical results
We employ the following values of model parameters that largely follow Cartea and Sánchez-Betancourt, (2022). For the price process, the value of the permanent price impact parameter is $\mathfrak{p}=10^{-3}$, the volatility is $\sigma^{S}=1$, and the initial value of the price is $S_{0}=100$. The parameter of the signal are as follows: the mean-reversion speed is $\kappa^{\alpha}=5$, the noise of the signal $\sigma^{\alpha}=1$, and the initial value is $\alpha_{0}=0$. The correlation between the noise in the price and the signal is zero ($\rho=0$). For the uninformed trading flow, we set $\kappa^{U}=15$, $\sigma^{U}=100$, and $\xi_{0}=0$. The temporary price impacts for the broker, informed, and uninformed trader are $\mathfrak{a}=2.1\times 10^{-3}$, $\mathfrak{b}=2\times 10^{-3}$, and $\mathfrak{c}=2\times 10^{-3}$, respectively.
In the informed trader’s learning equations, we set $\sigma^{B}=60$ and $\theta^{B}=10$. We assume that the initial estimate $\hat{\nu}_{0}$ is zero and we set the initial conditional variance $\mathbb{V}^{I}_{0}=0$. In the broker’s learning equations, we assume that she knows the values of the true parameters that describe the dynamics of $\alpha$, and similarly as in the informed trader’s framework, she sets $\alpha_{0}=0$ and $\mathbb{V}^{B}_{0}=0$. Both agents employ similar risk aversion parameters; we set $\beta_{0}^{I}=\beta_{0}^{B}=10^{-1}$, $\beta_{1}^{I}=\beta_{1}^{B}=10^{-3}$, $\rho_{0}^{I}=\rho_{0}^{B}=10^{-3}$, and $\rho_{1}^{I}=\rho_{1}^{B}=10^{-5}$. We run numerical simulations where we partition the time horizon $[0,T]$, with $T=1$, into $1,000$ equally-spaced intervals and we run $10,000$ simulations. [^8] The code is available at https://github.com/muhammadalifaqsha/broker_informed_noise_filtering_game.
Figure 1: The first three leading eigenvalues of $L(t)+L(t)^{\top}$. The fourth eigenvalue is zero.
Our parameters satisfy the conditions of Proposition 2 for existence and uniqueness of a solution. More precisely, Figure 1 shows the evolution of the first three eigenvalues of $L(t)+L(t)^{\top}$. Observe that the first three eigenvalues are negative, while the fourth eigenvalue is zero, as required in the proof of Proposition 2.
Figure 2: One sample path and $90\%$ confidence bands across simulations for the following processes. The midprice $S_{t}$, the signal $\alpha_{t}$, the broker’s trading speed $\nu_{t}$, the trading speed $\eta_{t}$ of the informed trader, the trading speed $\xi_{t}$ of the uninformed trader, the cash process $X^{I}$ of the informed trader, the cash process $X^{B}$ of the broker, the inventory $Q^{I}$ of the informed trader, and the inventory $Q^{B}$ of the broker.
Figure 2 shows one sample path for (i) the midprice $S_{t}$, (ii) the signal $\alpha_{t}$, (iii) the trading speeds of the broker, the informed, and the uninformed trader, (iv) the cash processes of the informed trader and the broker, and (vi) the inventory of both the informed trader and the broker. For each process, we plot the $90\%$ confidence bands across simulations.
Next, Figure 3 shows the estimates $\hat{\alpha}$ and $\hat{\nu}$. We plot the $90\%$ confidence bands across simulations. The bands around the filter are not the conditional variance of the filter, instead, they are the $90\%$ confidence bands for the filtered processes across simulations.
Figure 3: One realisation of filtering carried by both informed trader and broker.
Observe that the estimation accuracy of both filters is poor. This is ultimately due to the level of the permanent impact $\mathfrak{p}$ of the broker’s trading activity, and the level of the private signal $\alpha$ of the informed trader. The accuracy of our estimators are in line with practice; in financial markets, both agents face a low signal-to-noise ratio when using prices as a source of information to filter their competitor’s private information.
Lastly, Figure 4 shows the four components of the trading strategy of the broker. The main drivers are (i) inventory management (left panel), (ii) unwinding the uninformed order flow (third panel), and (iii) adjusting to the inventory level of the informed trader (fourth panel). In our simulation example, although the uninformed trading flow is profitable for the broker on average, the broker externalises part of it as seen from comparing third panel of Figure 4 and the path of $\xi$ in Figure 2. In particular, externalisation becomes more aggressive near the terminal time $T$ to reduce inventory risks. The rest of the flow is internalised.
Moreover, the last panel of Figure 4 and the last panel of Figure 2 show that the broker trades against the inventory level of the informed trader, in particular, she externalises a significant portion of the informed trading flow systematically. The second panel of Figure 4 shows the speculative component of the broker’s strategy which employs the estimated signal $\hat{\alpha}$. Observe that $\hat{\alpha}$ does not contribute to the broker’s strategy to the same extent as the other three components due to the poor estimation accuracy of the filter. Finally, the first panel of Figure 4 shows that the broker’s strategy unwinds the accumulated inventory (see last panel of Figure 2).
Figure 4: The four additive components of the broker’s strategy.
5.1   Performance and comparison against market practices
We compare the performance of the broker’s strategy in (4.77) with the following benchmarks. [^9] These benchmarks are those introduced in Cartea and Sánchez-Betancourt, (2022).
  1. (i)
    Benchmark 1: externalise the trading flow of the informed trader, internalise the trading flow of the uninformed trader, and unwind the broker’s inventory with TWAP. For this benchmark, the trading speed of the broker is:
    $\nu_{t}^{(1)}=\eta_{t}^{*}-\frac{Q_{t}^{B,(1)}}{T-t}\,.$ (5.1)
  2. (ii)
    Benchmark 2: internalise the trading flow of both the informed and uninformed traders, and unwind the broker’s inventory with TWAP. For this benchmark, the trading speed of the broker is:
    $\nu_{t}^{(2)}=-\frac{Q_{t}^{B,(2)}}{T-t}\,.$ (5.2)
  3. (iii)
    Benchmark 3: externalise the trading flow of both the informed and uninformed traders. For this benchmark, the trading speed of the broker is:
    $\nu_{t}^{(3)}=\eta_{t}^{*}+\xi_{t}\,.$ (5.3)
Below, we use $Q_{t}^{B,(i)}$, for $i\in\{1,2,3\}$, to denote the inventory of the broker when she follows strategy $\nu_{t}^{(i)}$.
We compute outperformance, defined as the additional trading performance from following the strategy (3.26), when compared to trading according to the benchmark strategies described above. Outperformance is expressed in dollars per million dollars traded. Mathematically, the outperformance compared to the $i$-th benchmark is expressed as
$\widehat{\text{Out}}^{(i)}=\frac{1}{N}\sum_{n=1}^{N}\left[\frac{X^{B,*}_{T,n}+Q^{B,*}_{T,n}\,S_{T,n}^{*}-\left(X^{B,(i)}_{T,n}+Q^{B,(i)}_{T,n}\,S_{T,n}^{(i)}\right)}{\int_{0}^{T}S_{u,n}^{(i)}\,\left(|\nu_{u,n}^{(i)}|+|\eta_{u,n}^{*}|+|\xi_{u,n}|\right)\,\mathrm{d}u}\right]\times 10^{6}\,,$ (5.4)
where $N=10,000$ simulations, $i\in\{1,2,3\}$ refers to the benchmark strategy, and the second sub-index is the simulation number. Table 1 shows the outperformance for each benchmark together with a $p$-value for the one-sided $t$-test with $H_{0}:\,\text{Out}^{(i)}=0$ against $H_{1}:\,\text{Out}^{(i)}>0$.
$i$ $\widehat{\text{Out}}^{(i)}$ (std) $p$-value
1 0 (570) $>0.5$
2 38 (354) $<0.001$
3 95 (695) $<0.001$
Table 1: Model outperformance with respect to the three benchmarks and their $p$-values from a one sided $t$-test ($H_{0}:\,\text{Out}^{(i)}=0$ and $H_{1}:\,\text{Out}^{(i)}>0$). The raw performance of the optimal strategy is 0.32 (2.81), and the performance for three benchmarks are 0.32 (2.39), 0.19 (2.71), and -0.04 (0.21).
Due to the low signal-to-noise ratio, the performance of the optimal strategy is indistinguishable from that of the first benchmark strategy which externalises the informed flow, internalises the uninformed flow, and unwinds inventory linearly. Note that we do not reject the null hypothesis (that the means are the same) in the $t$-test. In the next section, we study the performance of the broker’s strategy when she employs other sources of information to build her estimate of the private signal.
5.2   Performance with alternative estimators
Motivated by the poor performance of the broker when she estimates the signal from prices, we study the performance of the adaptive strategy (4.77) when the broker employs other estimates of the signal, including that described in Section 4.1.2.
5.2.1 Filtering directly from the informed trader’s speed
Here, we show that when the broker employs the estimate (4.28), which is constructed from the trading flow of the informed trader, she improves her trading performance significantly against the benchmarks described above. [^10] The performance reported here assumes no mispecification of inventory. Table 2 reports the outperformance. The performance we report here is a lower bound to the performance of the optimal strategy which uses the estimate (4.28) because the strategy we study here would be admissible in the adaptive control problem.
$i$ $\widehat{\text{Out}}^{(i)}$ (std) $p$-value
1 36 (117) $<0.001$
2 78 (389) $<0.001$
3 127 (518) $<0.001$
Table 2: Alternative estimation of $\alpha$ filtering directly from the informed trader’s speed. Model outperformance with respect to the three benchmarks and their $p$-values from a one sided $t$-test ($H_{0}:\,\text{Out}^{(i)}=0$ and $H_{1}:\,\text{Out}^{(i)}>0$).
We remark that the estimate (4.28) is only valid if (i) the informed trader carries out his trading exclusively with the broker and (ii) that he does not hide information or mislead the broker. If this is not the case, then the inventory of the informed trader, from the perspective of the broker, is mispecified, and the calculation of $\hat{\alpha}^{\text{alt}}_{t}$ above blows up at the terminal time $T$ as the limit in (4.16) blows up due to unaccounted inventory differences.
Figure 5: Left panel shows one sample path for the signal $\alpha$ and the new estimates $\hat{\alpha}$. The right panel shows the misspecified case for the informed trader’s inventory.
The left panel of Figure 5 shows the estimate assuming no mispecification in inventory, and the right panel shows a sample path when $Q^{I}_{0}\sim\mathcal{N}(0,1)$ and the broker assumes that $Q^{I}_{0}=0$. The mispecification of inventory has less impact on the accuracy of the estimate $\hat{\alpha}^{\text{alt}}$ in the beginning of the trading window because the informed trader trades less aggressively at this stage. However, as trading progresses, the accuracy deteriorates. Therefore, if the broker employs $\hat{\alpha}^{\text{alt}}$, then she must rely less on the signal less as the trading nears its close, since inventory mispecification increasingly contaminates the extracted signal at that stage.
Figure 6 shows the four components in the broker’s speed when she employs the new estimates $\hat{\alpha}^{\text{alt}}$. We observe that all components now contribute significantly to her trading strategy. In particular, the speculative component from exploiting $\hat{\alpha}^{\text{alt}}$ is more predominant than in Figure 4.
Figure 6: The four additive components of the broker’s strategy if she employs the new filtering strategy.
5.2.2 Naive estimate for the signal assuming near-zero values for $\hat{\nu}$
In practice, given that the permanent impact parameter $\mathfrak{p}$ of the broker’s trading activity in the market is small, the noise driving price movements obscures the filtered order flow of the broker. Thus, the contributions of the filtered trading speed $\hat{\nu}$ in the informed trader’s strategy are small enough so that they may be assumed to be zero. We illustrate this in Figure 7.
Figure 7: The components of the informed trader’s strategy.
The broker could obtain a better, near-perfect, estimate for $\alpha$. As shown in Figure 7, the contribution of $\hat{\nu}$ in the informed trader’s strategy is near-zero. Consequently, the broker could ignore those contributions (by assuming $f_{2}(t)\,\hat{\nu}_{t}=0$) and obtain the following alternative estimates for $\alpha$
$\hat{\alpha}^{\textup{naive}}_{t}:=\underbrace{\alpha_{t}+\frac{f_{2}(t)}{f_{1}(t)}\,\hat{\nu}_{t}}_{\approx\alpha_{t}}=\frac{\eta_{t}^{*}-f_{3}(t)\,Q_{t}^{I,*}}{f_{1}(t)}\,.$ (5.5)
The calculation for $\hat{\alpha}^{\textup{naive}}_{t}$ above always converges as $t\to T$ because
$\lim_{t\to T}\frac{f_{2}(t)}{f_{1}(t)}=\lim_{t\to T}\frac{f_{2}^{{}^{\prime}}(t)}{f_{1}^{{}^{\prime}}(t)}=\mathfrak{p}$
from L’Hôpital’s rule and the differential equations (3.20)–(3.21). Similar to the filter $\hat{\alpha}^{\text{alt}}$, we remark that this estimate is only valid if the informed trader carries out his trading exclusively with the broker. If this is not the case, then the inventory of the informed trader, from the perspective of the broker, is mispecified, and the calculation of $\hat{\alpha}_{t}$ above blows up at the terminal time $T$.
Figure 8 shows the differences between $\hat{\alpha}^{\textup{alt}}_{t}$ and $\hat{\alpha}^{\textup{naive}}_{t}$ with (right panel) and without (left panel) inventory mispecification. The differences in the left panel are negligible for practical purposes (the $y$-axis is in units of $10^{-5}$), and the differences in the right panel are in the same order of magnitude as the mispecification. As expected, the differences blow up when $t\uparrow T$.
Figure 8: Both panels show the differences between $\hat{\alpha}^{\textup{alt}}_{t}$ and $\hat{\alpha}^{\textup{naive}}_{t}$ with a 90% confidence band over simulations. Left panel is for no mispecification of inventory. The right panel has mispecification of inventory, that is, $Q^{I}_{0}\sim\mathcal{N}(0,1)$ and the broker assumes that $Q^{I}_{0}=0$.
Finally, we compute the outperformance over the benchmarks when the broker employs the estimate $\hat{\alpha}^{\textup{naive}}$ of $\alpha$ above, and assuming no mispecification of inventory. As expected, we find that the outperformances are the same as those in Table 2. That is, if we compare the performance obtained using the two filtered signals in this section, a $t$-test does not find evidence to reject the null hypothesis that their mean performances are the same.
5.3   Second-order effects
In the problem formulation above, both agents act strategically, though the informed trader’s strategy is only strategic to a first order. Specifically, the informed trader does not account for the possibility that the broker may filter his private signal. Conversely, the broker operates under the assumption that the informed trader will incorporate her trading activity to enhance trading performance. Here, we explore how to incorporate this second-order effect into the informed trader’s problem and evaluate the value of the broker’s second-order effect.
5.3.1 Informed trader
From the perspective of the informed trader, a second-order effect can be incorporated by assuming that the broker’s trading follows the model
$\nu_{t}=c_{t}\,\eta_{t}+\underline{\nu}_{t}\,,$ (5.6)
where $c_{t}$ is a deterministic function that reflects how the broker externalises his trading flow, and $\underline{\nu}_{t}$ is the remainder of the broker’s trading flow. In this setup, the informed trader assumes that $\underline{\nu}$ is exogenous and mean reverting, as in the initial model. Given this assumption, and similar to the approach above, the informed trader can filter $\underline{\nu_{B}}$ to compute $\hat{\underline{\nu}}$, and then formulate an adaptive control problem. This control problem would be similar to that in (3.10), but where one changes $\hat{\nu}$ to $\underline{\hat{\nu}}$, and adds a term of the form $\mathfrak{p}\,c_{t}\,\eta_{t}\,\partial_{s}H^{I}$. While we omit the detailed calculations here, the intuition is that the informed trader benefits in this modified model, as he effectively “observes” more of the broker’s trading speed, which enhances his strategic advantage.
Another approach is for the informed trader to assume that the broker’s trading follows the model
$\nu_{t}=c_{t}^{1}\,\hat{\alpha}_{t}+c_{t}^{2}\,Q^{I}_{t}+\underline{\nu}_{t}\,,$ (5.7)
where $\hat{\alpha}_{t}$ is the broker’s estimate of the signal, which the informed trader cannot construct independently. Similar to the broker’s problem in Section 4, the informed trader could substitute $\hat{\alpha}$ with her true signal and filter $\underline{\nu}_{t}$, assuming the remaining trading flow of the broker is mean reverting and exogenous. Mathematically, the price process would follow the dynamics
$\mathrm{d}S_{t}=\left[\mathfrak{p}\,\hat{\underline{\nu}}_{t}+(1+\mathfrak{p}\,c_{t}^{1})\,\alpha_{t}+\mathfrak{p}\,c_{t}^{2}\,Q^{I}_{t}\right]\,\mathrm{d}t+\sigma^{S}\,\mathrm{d}\tilde{W}^{S,I}_{t}\,,$ (5.8)
where $\hat{\underline{\nu}}_{t}$ is the filter of ${\underline{\nu}}_{t}$.
5.3.2 Broker
To study the impact of the second-order effects in the strategy of the broker, we introduce the following modification to the initial broker’s assumption (4.2) regarding the informed trader’s strategy:
$\eta^{*}_{t}=f_{1}(t)\,\hat{\alpha}_{t}+c\,f_{2}(t)\,\nu_{t}+f_{3}(t)\,Q^{I*}_{t}\,,$ (5.9)
where the new parameter $c$ is employed to assess the degree to which the broker believes she can influence the behaviour of the informed trader.
Based on the model assumption (5.9), the broker computes her optimal speed $\nu^{*,c}$. Note that the case $c=1$ corresponds to the initial model described in Section 4. Our aim is now to examine how $\nu^{*,c}$ deviates from the optimal strategy in the original model where $c=1$. To study the manipulation effects, we reduce the influence of the risk terms on the strategies of both the broker and the informed trader by setting $\beta_{0}^{I}=\beta_{0}^{B}=10^{-5}$.
Figure 9: Scatter plot of $\nu^{*,c=0}$, $\nu^{*,c=1}$, and $\nu^{*,c=1}-\nu^{*,c=0}$ against the inventory of the broker for $c=0$ and $c=1$.
The broker’s trading speed is a component in the informed trader’s strategy (5.9). Thus, in our original model, the assumption (5.9) may be interpreted as a certain “belief” of the broker in her ability to influence the informed trader’s strategy. Figure 9 shows that the broker uses this component to her advantage. More precisely, the broker will, when profitable, intentionally mislead the informed trader into trading in the opposite direction of her inventory. The aim is to offload her inventory to her client rather than unwinding it in the lit market herself.
5.4   Mispecification of learning parameters
Here, we examine how mispecifications in model parameter values affect the outperformance metrics reported above. Using the values reported in Table 1 as a baseline, we stress each parameter by increasing and decreasing its value by $50\%$, and report the results in Table 3.
$i$ Base $\kappa^{\alpha}$ (+50%) $\kappa^{\alpha}$ (-50%) $\sigma^{\alpha}$ (+50%) $\sigma^{\alpha}$ (-50%)
1 0 (570) -11 (572) -12 (569) -12 (569) -12 (574)
2 38 (354)* 38 (358)* 37 (364)* 37 (365)* 38 (359)*
3 95 (695)* 88 (695)* 87 (693)* 87 (693)* 88 (696)*
$i$ Base $\theta^{B}$ (+50%) $\theta^{B}$ (-50%) $\sigma^{B}$ (+50%) $\sigma^{B}$ (-50%)
1 0 (570) -11 (569) -12 (573) -12 (573) -11 (566)
2 38 (354)* 38 (356)* 37 (362)* 37 (362)* 38 (351)*
3 95 (695)* 88 (693)* 87 (696)* 87 (696)* 88 (691)*
Table 3: Outperformance of model to the three benchmarks when the broker mispecified her learning parameters. Significant cases are marked with *.
We observe that, even with parameter mispecification, the broker’s optimal strategy continues to outperform benchmarks 2 and 3, with outperformance levels remaining approximately the same as in the baseline. However, the broker’s outperformance over the first benchmark now turns negative; note that these values are not statistically significant.
5.5   The role of externalisation and stressing model parameters
In this section, we examine the role of externalisation in the trading strategy of the broker. Additionally, we analyse the impact of changes in model parameters on the externalisation rate. Specifically, we focus on (i) the parameters in the dynamics of the signal and (ii) the parameters of the informed trader’s assumption over the broker’s trading strategy.
Remark 3.
If the broker is correct in her estimate of $\alpha$ and if the informed trader only conducts his trade with the broker (the broker has a good estimate of $Q^{I}$), then, the broker’s “externalisation” that happens in advance of the trades taking place, might be catalogued as “frontrunning”. For example, the broker might buy (resp. sell) in the lit market in advance of the informed trader sending his buy (resp. sell) instruction. However, we remark that these actions are taken based on a probabilistic model as opposed to being taken based on factual knowledge of what the client is going to do next.
Figure 10 shows the “effective” externalisation quotient for stressed values of model parameters. To derive this quotient, observe that $\nu^{*}_{t}=f^{B}_{\alpha}(t)\,\hat{\alpha}_{t}+\epsilon_{t}$ and similarly, that $\eta^{*}_{t}=f^{I}_{\alpha}(t)\,\alpha_{t}+\varepsilon_{t}$ for some $\{\epsilon_{t},\varepsilon_{t}\}$. Thus, a straightforward calculation shows that if $\alpha_{t}\approx\hat{\alpha}_{t}$, then $\nu^{*}_{t}=f^{B}_{\alpha}(t)/f^{I}_{\alpha}(t)\,\eta^{*}_{t}+\upsilon_{t}$. In this setup, the quotient $f^{B}_{\alpha}(t)/f^{I}_{\alpha}(t)$ represents the effective externalisation rate at which the broker offloads the trading of the informed trader. This quotient is representative of the externalisation rate in practice when the broker has a good estimate of the private signal.
Figure 10: Quotient $f^{B}_{\alpha}/f^{I}_{\alpha}$ that represents the effective externalisation followed by the broker. This quotient accounts for the speculation by the broker according to her signal estimates.
Figure 10 shows that the mean-reversion parameter $\kappa^{\alpha}$ in the signal dynamics influences the effective externalisation rate significantly. Large values of $\kappa^{\alpha}$ result in a signal that diffuses less. As such, the broker trusts her estimates more and increases her speculation part, thus she externalises more. As the signal diffuses less (from zero), its role diminishes in the strategies of both the broker and informed trader; see Figure 11. Visually, the signal appears to impact both the broker and informed trader’s strategies similarly. However, the broker’s sensitivity to the signal decreases at a faster rate when $\kappa^{\alpha}$ is low (i.e. when the signal diffuses more). This can be attributed to the broker’s growing uncertainty in her estimate of the signal due to accumulated errors. When $\kappa^{\alpha}$ is high (i.e. when the signal diffuses less), the broker’s sensitivity to the signal decreases at a more equal rate to the trader. Consequently, the broker increases her externalisation rate without concern for the risks associated with signal misspecification.
Figure 11: The coefficients entering the estimated signal component in the broker (left panel) and informed trader’s strategy (right panel).
[…]
Figure 12: The left panel is for when the broker uses the original estimates $\hat{\alpha}$, and the right panel is for the naive filter $\hat{\alpha}^{\text{naive}}$. In both panels we show the median trajectory when stressing the values of model parameters.
Figure 12 illustrates that, with the original filtering framework, the broker’s externalisation rate is centred around zero. In contrast, the median externalisation in the naive case (right panel) is significantly higher, ranging between $70$ and $90\%$. This discrepancy is due to the difference in signal estimation quality between $\hat{\alpha}$ and $\hat{\alpha}^{\text{naive}}$. The ordering of the lines when stressing $\kappa^{\alpha}$ in the right panel of Figure 12 is similar to those in Figure 10.
Appendix A Proof of Proposition 1
We need the following lemmas to derive the filtering equations from the dynamics of the informed trader’s speed. Recall that the process $Z$ (which is observable) follows the equation
$\begin{split}\mathrm{d}Z_{t}&=\mathfrak{G}_{7}(t)\,\alpha_{t}\,\mathrm{d}t+\mathrm{d}\tilde{W}_{t}\end{split}$ (A.1)
where the signal $\alpha$ is latent and $\tilde{W}$ is given in the lemma below.
Lemma 2.
The process below
$\mathrm{d}\tilde{W}_{t}:=\frac{1}{\mathfrak{G}_{5}(t)}\,\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]$ (A.2)
is a $(\mathbb{P},(\mathcal{F}_{t})_{t})$-Brownian motion. Furthermore, the process below (called the innovation process)
$I_{t}:=Z_{t}-\int_{0}^{t}\mathbb{E}\left[\mathfrak{G}_{7}(s)\,\alpha_{s}\,\Big{|}\,\mathcal{Z}_{s}\right]\,\mathrm{d}s$ (A.3)
is a $(\mathbb{P},(\mathcal{Z}_{t}))$-Brownian motion.
Proof.
The first part of the lemma is obvious as the Ito integrand are bounded by $1$ and the quadratic variation is $\mathrm{d}t$. Next, we will prove that it is a $(\mathbb{P},(\mathcal{Z}_{t}))$-martingale. Let $0\leq s\leq t\leq T$. Then
$\begin{split}\mathbb{E}\left[I_{t}-I_{s}\,|\,\mathcal{Z}_{s}\right]&=\mathbb{E}\left[Z_{t}-Z_{s}-\int_{s}^{t}\mathbb{E}\left[\mathfrak{G}_{7}(u)\,\alpha_{u}\,\Big{|}\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,\Big{|}\,\mathcal{Z}_{s}\right]\\ &=\mathbb{E}\left[Z_{t}-Z_{s}\,|\,\mathcal{Z}_{s}\right]-\mathbb{E}\left[\int_{s}^{t}\mathfrak{G}_{7}(u)\,\alpha_{u}\,\mathrm{d}u\,\Big{|}\,\mathcal{Z}_{s}\right]\,.\end{split}$ (A.4)
Note that the diffusion terms in (A.2) are bounded by $1$, so that
$\mathbb{E}\left[Z_{t}-Z_{s}\,|\,\mathcal{F}_{s}\right]=\mathbb{E}\left[\int_{s}^{t}\mathfrak{G}_{7}(u)\,\alpha_{u}\,\mathrm{d}u\Big{|}\,\mathcal{F}_{s}\right]\,.$
But $\mathcal{Z}_{s}\subset\mathcal{F}_{s}$, so that $\mathbb{E}[Z_{t}-Z_{s}\,|\mathcal{Z}_{s}]=\mathbb{E}[\mathbb{E}[Z_{t}-Z_{s}\,|\,\mathcal{F}_{s}]\,|\,\mathcal{Z}_{s}]$, and we obtain (A.4) equals zero.
Next, it can be easily calculated that $\langle I\rangle_{t}=t$. By Levy Characterisation, $I$ is a $(\mathbb{P},(\mathcal{Z}_{t})_{t})$-Brownian motion. ∎
Lemma 3.
The processes
$M_{t}$ $=\mathbb{E}\left[-\int_{0}^{t}\kappa^{\alpha}\,\alpha_{u}\,\mathrm{d}u+\sigma^{\alpha}\,W_{t}^{\alpha}\,\big{|}\,\mathcal{Z}_{t}\right]+\int_{0}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,,$ (A.5)
$N_{t}$ $=\mathbb{E}\left[-\int_{0}^{t}\kappa^{\alpha}\,\alpha_{u}^{2}\,\mathrm{d}u+\int_{0}^{t}\sigma^{\alpha}\,\alpha_{u}\,\mathrm{d}W_{u}^{\alpha}\,\big{|}\,\mathcal{Z}_{t}\right]+\int_{0}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,$ (A.6)
are $(\mathbb{P},(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}})$-martingales.
Proof.
Observe that for $s\leq t$, and the fact that $W^{\alpha}$ is a $(\mathbb{P},(\mathcal{F}_{t})_{t\in{\mathfrak{T}}})$-Brownian motion, we may obtain the following by using the tower property.
$\mathbb{E}\left[M_{t}-M_{s}\,\big{|}\,\mathcal{Z}_{s}\right]$ $=\mathbb{E}\left[-\int_{s}^{t}\kappa^{\alpha}\,\alpha_{u}\,\mathrm{d}u\,\big{|}\,\mathcal{Z}_{s}\right]+\mathbb{E}\left[\int_{s}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,\big{|}\,\mathcal{Z}_{s}\right]=0\,.$ (A.7)
Additionally, $\mathbb{E}\left[\int_{0}^{T}\alpha_{u}^{2}\,\mathrm{d}u\right]<\infty$ because $\mathbb{E}\left[\alpha_{u}^{2}\right]=\frac{(\sigma^{\alpha})^{2}}{2\,\kappa^{\alpha}}\exp(-2\,\kappa^{\alpha}\,u)$. Thus, $\left(\int_{0}^{t}\alpha_{u}\,\mathrm{d}W_{u}^{\alpha}\right)_{t\in{\mathfrak{T}}}$ is a $(\mathbb{P},(\mathcal{F}_{t})_{t\in{\mathfrak{T}}})$-martingale. Using the same trick as before,
$\mathbb{E}\left[N_{t}-N_{s}\,\big{|}\,\mathcal{Z}_{s}\right]$ $=\mathbb{E}\left[-\int_{s}^{t}\kappa^{\alpha}\,\alpha_{u}^{2}\,\mathrm{d}u\,\big{|}\,\mathcal{Z}_{s}\right]+\mathbb{E}\left[\int_{s}^{t}\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,|\,\mathcal{Z}_{u}\right]\,\mathrm{d}u\,\big{|}\,\mathcal{Z}_{s}\right]=0\,.$ (A.8)
Proof for Proposition 1.
From Lemma 3, $M$ is a $(\mathbb{P},(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}})$-martingale. Because and $I$ is a $(\mathbb{P},(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}})$-Brownian motion, by the martingale representation theorem, there exists a $(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}}$-adapted process $r^{M}$ such that $M_{t}=\int_{0}^{t}r_{u}^{M}\,\mathrm{d}I_{u}$. We compute
$\hat{\alpha}^{\text{alt}}_{t}=\mathbb{E}\left[-\int_{0}^{t}\kappa^{\alpha}\,\alpha_{u}\,\mathrm{d}u+\sigma^{\alpha}\,W_{t}^{\alpha}\,\big{|}\,\mathcal{Z}_{t}\right]=M_{t}-\int_{0}^{t}\kappa^{\alpha}\,\hat{\alpha}^{\text{alt}}_{u}\,\mathrm{d}u=\int_{0}^{t}r_{u}^{M}\,\mathrm{d}I_{u}\ -\int_{0}^{t}\kappa^{\alpha}\,\hat{\alpha}^{\text{alt}}_{u}\,\mathrm{d}u\,.$ (A.9)
By applying Ito’s formula to $(\hat{\alpha}^{\text{alt}}_{t})^{2}$, we obtain
$\mathrm{d}(\hat{\alpha}^{\text{alt}}_{t})^{2}=\left[-2\,\kappa^{\alpha}\,(\hat{\alpha}^{\text{alt}}_{t})^{2}+(r_{t}^{M})^{2}\right]\,\mathrm{d}t+2\,r_{t}^{M}\,\hat{\alpha}^{\text{alt}}_{t}\,\mathrm{d}I_{t}\,.$ (A.10)
Similarly for $N_{t}$, there exists a $(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}}$-adapted process $r^{N}$ such that $N_{t}=\int_{0}^{t}r_{u}^{N}\,\mathrm{d}I_{u}$. Now, we apply Ito’s formula to $\alpha_{t}^{2}$ and take its conditional expectation with respect to $\mathcal{Z}_{t}$.
$\begin{split}\mathbb{E}\left[\alpha_{t}^{2}\,\big{|}\,\mathcal{Z}_{t}\right]&=\int_{0}^{t}\Big{\{}(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,\big{|}\,\mathcal{Z}_{t}\right]\Big{\}}\,\mathrm{d}u+\sigma^{\alpha}\,\mathbb{E}\left[\int_{0}^{t}2\,\alpha_{u}\,\mathrm{d}W_{u}^{\alpha}\,\big{|}\,\mathcal{Z}_{t}\right]\\ &=\int_{0}^{t}2\,r_{u}^{N}\,\mathrm{d}I_{u}+\int_{0}^{t}\Big{\{}(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{E}\left[\alpha_{u}^{2}\,\big{|}\,\mathcal{Z}_{u}\right]\Big{\}}\,\mathrm{d}u\,.\end{split}$ (A.11)
From Lemma 2, $\tilde{W}$ is a $(\mathbb{P},(\mathcal{F}_{t})_{t\in{\mathfrak{T}}})$-Brownian motion. Because the signal $\alpha$ is a $\mathbb{P}$-Gaussian process, then $Z$ is also a $\mathbb{P}$-Gaussian process. Thus, the projection $\alpha_{t}$ into $\mathcal{Z}_{t}$ is also Gaussian. Thus, $\alpha_{t}-\hat{\alpha}^{\text{alt}}_{t}$ is independent of $\mathcal{Z}_{t}$, so that $\mathbb{V}^{\text{alt},B}_{t}$ must be a deterministic quantity. As such, we should have $r_{t}^{M}\,\hat{\alpha}^{\text{alt}}_{t}=r_{u}^{N}$, and
$\mathrm{d}\mathbb{V}^{\text{alt},B}_{t}=\Big{\{}(\sigma^{\alpha})^{2}-2\,\kappa^{\alpha}\,\mathbb{V}^{\text{alt},B}_{t}-(r_{t}^{M})^{2}\Big{\}}\,\mathrm{d}t\,.$ (A.12)
[…]
$\mathrm{d}\left(\epsilon\,\alpha\right)_{t}$ $=\epsilon_{t}\,\mathrm{d}\alpha_{t}+\alpha_{t}\,\mathrm{d}\epsilon_{t}+\mathrm{d}\langle\epsilon,\alpha\rangle_{t}$ (A.13)
$=-\kappa^{\alpha}\,\epsilon_{t}\,\alpha_{t}\,\mathrm{d}t+\sigma^{\alpha}\,\epsilon_{t}\,\mathrm{d}W_{t}^{\alpha}+i\,\epsilon_{t}\,\alpha_{t}\,r(t)\,\mathrm{d}Z_{t}+i\,\sigma^{\alpha}\,\epsilon_{t}\,r(t)\,\mathfrak{K}(t)\,\mathrm{d}t\,,$ (A.14)
$\mathrm{d}\left(\epsilon\,\hat{\alpha}^{\text{alt}}\right)_{t}$ $=\epsilon_{t}\,\mathrm{d}\hat{\alpha}^{\text{alt}}_{t}+\hat{\alpha_{t}}^{\text{alt}}\,\mathrm{d}\epsilon_{t}+\mathrm{d}\langle\epsilon,\hat{\alpha}^{\text{alt}}\rangle_{t}$ (A.15)
$=-\kappa^{\alpha}\,\epsilon_{t}\,\hat{\alpha}^{\text{alt}}_{t}\,\mathrm{d}t+\epsilon_{t}\,r^{M}_{t}\,\mathrm{d}I_{t}+i\,\epsilon_{t}\,r(t)\,\hat{\alpha}^{\text{alt}}_{t}\,\mathrm{d}Z_{t}+i\,\epsilon_{t}\,r(t)\,r_{t}^{M}\,\mathrm{d}t\,.$ (A.16)
Note that by applying the tower property, $\mathbb{E}\left[\epsilon_{t}(\hat{\alpha}^{\text{alt}}_{t}-\alpha_{t})\right]=0$. By subtracting (A.14) from (A.16) and taking the total expectation, we have
[…]
We know that $\mathbb{V}^{\text{alt},B}_{t}\leq\text{Var}(\alpha_{t})=\frac{(\sigma^{\alpha})^{2}}{2\,\kappa^{\alpha}}\,\exp\left(-2\,\kappa^{\alpha}\,t\right)$, so that $\mathbb{E}\left[\int_{0}^{T}\left(\hat{\alpha}^{\text{alt}}_{u}-\alpha_{u}\right)^{2}\,\mathrm{d}u\right]<\infty$. Thus, the rightmost term in (A.17) equals zero. We obtain
$0=\int_{0}^{t}i\,r(u)\,\mathbb{E}\left[\epsilon_{u}\,\left\{r_{u}^{M}-\sigma^{\alpha}\,\mathfrak{K}(u)+(\hat{\alpha}^{\text{alt}}_{u}-\alpha_{u})\,\mathfrak{G}_{7}(u)\,\alpha_{u}\right\}\right]\,\mathrm{d}u,\quad\forall{t\in{\mathfrak{T}}},$
and consequently,
$\mathbb{E}\left[\epsilon_{t}\,\left\{r_{t}^{M}-\sigma^{\alpha}\,\mathfrak{K}(t)+(\hat{\alpha}^{\text{alt}}_{t}-\alpha_{t})\,\mathfrak{G}_{7}(t)\,\alpha_{t}\right\}\right]=0,\quad\forall{t\in{\mathfrak{T}}}.$
This equation applies for all $r\in L^{\infty}([0,T];\mathbb{R})$. Using Lemma B.39 in Bain and Crisan, (2009), we have $\mathbb{E}\left[r_{t}^{M}-\sigma^{\alpha}\,\mathfrak{K}(t)+(\hat{\alpha}_{t}-\alpha^{\text{alt}}_{t})\,\mathfrak{G}_{7}(t)\,\alpha_{t}\,\big{|}\,\mathcal{Z}_{t}\right]=0$ almost surely. Thus, we obtain
$r_{t}^{M}=\sigma^{\alpha}\,\mathfrak{K}(t)+\mathfrak{G}_{7}(t)\,\mathbb{V}^{\text{alt},B}_{t}\,.$ (A.18)
Substituting this into (A.9) and (A.12), we obtain (4.27) and (4.28) in Proposition 1. ∎